### v2

# Bessel potentials and Green functions on pseudo-Euclidean spaces

Abstract
We review properties of Bessel potentials, that is, inverse Fourier transforms of (regularizations of) $(m^{2}+p^{2})^{-\frac{\mu}{2}}$ on a pseudoEuclidean space with signature $(q,d-q)$. We are mostly interested in the Lorentzian signature $(1,d-1)$, and the case $\mu=2$, related to the Klein-Gordon equation $(-\Box+m^{2})f=0$. We analyze properties of various “propagators”, which play an important role in Quantum Field Theory, such as the retarded/advanced propagators or Feynman/antiFeynman propagators. We consistently use hypergeometric functions instead of Bessel functions, which makes most formulas much more transparent. We pay attention to distributional properties of various Bessel potentials. We include in our analysis the “tachyonic case”, corresponding to the “wrong” sign in the Klein-Gordon equation.
Keywords: Bessel potential, Riesz potential, Klein-Gordon equation, Minkowski space.
1 Introduction
Let us start with the Bessel potentials on the Euclidean space ${\mathbb{R}}^{d}$. Let ${\rm Re}\mu>0$ and $m\geq 0$. If $m=0$ we will usually additionally assume that $d>{\rm Re}\mu$. Consider the function
$G_{\mu,m}(x)=\int\frac{{\rm e}^{{\rm i}px}}{(m^{2}+p^{2})^{\frac{\mu}{2}}}\frac{{\rm d}p}{(2\pi)^{d}}$ (1.1)
on the Euclidean space ${\mathbb{R}}^{d}$. Note that that $G_{\mu,m}(x-y)$ can be interpreted as the integral kernel of the operator $(m^{2}-\Delta)^{-\frac{\mu}{2}}$.
We have
$G_{\mu,m}(x)=m^{d-\mu}G_{\mu,1}(mx),$ (1.2)
so the case $m>0$ reduces to $m=1$. $G_{\mu,1}(x)$ can be expressed in terms of the Macdonald function, one of solutions of the modified Bessel equation. Therefore, $G_{\mu,1}(x)$ is often called the Bessel potential of order $\mu$. The function $G_{\mu,0}(x)$ is called the Riesz potential of order $\mu$.
It is remarkable that the theory of Bessel potentials is very similar for all $\mu>0$. However, the case $\mu=2$ is probably the most important. In this case we will usually omit $\mu$ from the notation, setting $G_{m}(x):=G_{2,m}(x)$, and obtaining the Green function of the inhomogeneous Helmholtz equation
$(-\Delta+m^{2})g(x)=f(x).$ (1.3)
In other words,
$(-\Delta+m^{2})G_{m}(x)=\delta(x),$ (1.4)
Note that in dimension $d=3$ we have $G_{m}(x)=\frac{{\rm e}^{-m|x|}}{4\pi|x|}$. Thus for $m>0$ it coincides with the Yukawa potential and for $m=0$ with the Coulomb potential.
Suppose now ${\mathbb{R}}^{q,d-q}$ is the pseudo-Euclidean space of signature $(q,d-q)$. In other words, as a set it is ${\mathbb{R}}^{d}$ with the scalar product for $x,y\in{\mathbb{R}}^{q,p}$ given by
$xy=-x_{1}y_{1}\cdots-x_{q}y_{q}+x_{q+1}y_{q+1}+\cdots x_{d}y_{d}.$ (1.5)
The definition (1.1) is usually no longer correct for $m^{2}\in{\mathbb{R}}$, since $\frac{1}{(m^{2}+p^{2})^{\frac{\mu}{2}}}$ may fail to be locally integrable, and hence may not define a tempered distribution. It still works for complex non-real $m^{2}$. A possible pair of generalizations of (1.1) to $m^{2}$ real is the pair of functions, which correspond to the limits from above and below:
$G_{\mu,m}^{\mathrm{F}}(x)=\int\frac{{\rm e}^{{\rm i}px}}{(m^{2}+p^{2}-{\rm i}0)^{\frac{\mu}{2}}}\frac{{\rm d}p}{(2\pi)^{d}},$ (1.6)
$G_{\mu,m}^{\overline{\mathrm{F}}}(x)=\int\frac{{\rm e}^{{\rm i}px}}{(m^{2}+p^{2}+{\rm i}0)^{\frac{\mu}{2}}}\frac{{\rm d}p}{(2\pi)^{d}}.$ (1.7)
(1.6) and (1.7) have an obvious interpretation as boundary values of integral kernels of appropriate functions of the pseudoLaplacian
$\Box:=-\partial_{1}^{2}\cdots-\partial_{q}^{2}+\partial_{q+1}^{2}+\partial_{d}^{2}.$ (1.8)
Again, the case $m>0$ reduces to $m=1$. $G_{\mu,m}^{\mathrm{F}/\overline{\mathrm{F}}}(x)$ can be expressed by Macdonald and Hankel functions. (The Hankel functions are special functions solving the standard Bessel equation.)
The symbols $\mathrm{F}$ and $\overline{\mathrm{F}}$ are motivated by the special case of Green functions in the Lorentzian case. $G_{2,m}^{\mathrm{F}/\overline{\mathrm{F}}}(x)$ coincide then with the Feynman, resp. the anti-Feynman propagators, which play an important role in Quantum Field Theory, as we explain below.
In our paper we will discuss all signatures, including the Euclidean $(0,d)$ and anti-Euclidean $(d,0)$. However, we are mostly interested in the Lorentzian signature. The Lorentzian signature comes in two varieties: “mostly pluses” $(1,d-1)$ and “mostly minuses” $(d-1,1)$. We will treat the former as the standard one.
The Lorentzian case is especially interesting and rich. This is related to the fact that the Minkowski space ${\mathbb{R}}^{1,d-1}$ can be equipped with a causal structure and the set $p^{2}+m^{2}=0$ has two connected components. Therefore, besides $G_{\mu,m}^{\mathrm{F}/\overline{\mathrm{F}}}$, we can introduce the distributions
$G_{\mu,m}^{\vee}(x)=\int\frac{{\rm e}^{{\rm i}px}}{(m^{2}+p^{2}-{\rm i}0{\rm sgn}p^{0})^{\frac{\mu}{2}}}\frac{{\rm d}p}{(2\pi)^{d}},$ (1.9)
$G_{\mu,m}^{\wedge}(x)=\int\frac{{\rm e}^{{\rm i}px}}{(m^{2}+p^{2}+{\rm i}0{\rm sgn}p^{0})^{\frac{\mu}{2}}}\frac{{\rm d}p}{(2\pi)^{d}},$ (1.10)
which are invariant wrt orthochronous Lorentz transformations. Remarkably, $G_{\mu,m}^{\vee/\wedge}$ is supported in the forward, resp. backward cone. Therefore, $G_{\mu,m}^{\vee}$ is called the forward (or retarded), and $G_{\mu,m}^{\wedge}$ the backward (or advanced) Bessel potential.
In the Lorentzian case, the pseudo-Laplacian is usually called the d’Alembertian
$\Box:=-\partial_{0}^{2}+\partial_{1}^{2}+\cdots+\partial_{d-1}^{2},$ (1.11)
and $-\Box+m^{2}$ is called the Klein-Gordon operator. By a Green function of the (inhomogeneous) Klein-Gordon equation
$(-\Box+m^{2})f(x)=g(x).$ (1.12)
we will mean a distribution $G^{\bullet}(x)$ satisfying
$(-\Box+m^{2})G^{\bullet}(x)=\delta(x).$ (1.13)
The Klein-Gordon equation possesses many Green functions. Among them, we have the Feynman and antiFeynman Green functions given by the formulas (1.6) and (1.7) with $\mu=2$. Another distinguished pair consists of the retarded (or forward) Green function and the advanced (or backward) Green function, defined by demanding that their support is contained in the forward, resp. backward cone. For $m^{2}\geq 0$ the retarded Green function is given by (1.9) and the advanced Green function by (1.10) with $\mu=2$.
The Feynman, anti-Feynman, forward, and backward Green functions of the Klein-Gordon equation have important applications in physics, especially in classical and quantum field theory. The forward and backward Green functions can be used to express the Cauchy problem. The Feynman, resp. anti-Feynman Green functions express the time-ordered, resp. anti-time-ordered vacuum expectation values of fields in quantum field theory. Importantly, they satisfy the identity
$G_{m}^{\mathrm{F}}+G_{m}^{\overline{\mathrm{F}}}=G_{m}^{\vee}+G_{m}^{\wedge}.$ (1.14)
In our paper, we also consider the Lorentzian case with the “wrong sign of $m^{2}$”. This case corresponds to the tachyonic Klein-Gordon equation
$(-\Box-m^{2})f(x)=g(x).$ (1.15)
Remarkably, all four basic Green functions, Feynman $G_{m}^{\mathrm{F}}$, anti-Feynman $G_{m}^{\overline{\mathrm{F}}}$, forward $G_{m}^{\vee}$, and backward $G_{m}^{\wedge}$, can be defined in the tachyonic case. For the Feynman and anti-Feynman Green functions we can still use the formulas (1.6) and (1.7), where $m^{2}$ is replaced with $-m^{2}$. Their interpretation in terms of the vacuum expectation values is however lost, since the tachyonic theory has no vacuum state. (In particular, in the tachyonic case we do not have a counterpart of the positive/negative frequency Green functions (5.40)). The forward and backward Green functions are defined by their support properties. For them we cannot use the formulas (1.9) and (1.10). In fact, the set $p^{2}-m^{2}=0$ is now connected, and cutting it with ${\rm sgn}p^{0}$ is no longer invariant. Nevertheless, one can use the analytic continuation in $m$ to uniquely define Green functions with correct support properties also in the tachyonic case. We point out that the identity (1.14) is no longer true in the tachyonic case.
The difference of two Green functions is a solution of the homogeneous Helmholtz/Klein-Gordon equation. Certain distinguished solutions are important for physics applications. In the Lorentzian case, we have the Pauli-Jordan propagator; for $m^{2}\geq 0$ also the positive frequency and the negative frequency two-point functions. We illustrate applications of distinguished solutions to the Helmholtz/Klein-Gordon equation by computing averages of plane waves over the sphere (in the Euclidean case), as well as over the hyperbolic and de Sitter space (in the Lorentzian case).
Let us say a few words about the history of Bessel potentials. The name Bessel potentials was introduced in the 60s by Aronszajn and Smith, who studied them in the Euclidean case in [1]. Around the same time, they were also investigated by Calderon [2]. Bessel potentials are frequently viewed in the literature as smoothed versions of Riesz potentials (see, for example, [3] where they are defined using the integral formula (2.5)). They are often used to define Bessel potential spaces that generalize standard Sobolev spaces (see [4]), and the idea to use Bessel kernels is due to Deny [5]. For a comprehensive treatment of (Euclidean) Bessel potentials, we refer the reader to [1], where many properties of Bessel potentials are exhaustively studied.
The Lorentzian versions of Bessel potentials, typically in dimension 1+3, often appear in the literature on Quantum Field Theory. They are ingredients of formulas for scattering amplitudes based on Feynman diagrams and on the Epstein-Glaser approach [6, 7]. The famous textbooks by Björken-Drell [8] and by Bogoliubov– Shirkov [9] contain appendices devoted to distinguished Green functions and solutions of the Klein-Gordon equation in the physical dimension 1+3. They carry various names. For instance, often the term Green function is replaced by propagator, etc.
Formulas for Bessel potentials in various signatures are known and are available in collections of integrals such as [10] and [11]. In chapter III.2 of [12] one can find Fourier transforms of powers of quadratic forms with any signature, including the formula (4.5) of the general case studied in this paper. Although there exists a large literature about Bessel potentials, our presentation contains several new points, which we have not seen in the literature and believe are important.
The first new point involves the special functions that we use. Various kinds of the Bessel equation can be reduced to equation
$(z{\rm\partial}_{z}^{2}+(\alpha+1){\rm\partial}_{z}-1)v(z)=0,$ (1.16)
which can be called the ${}_{0}F_{1}$ hypergeometric equation. Equation (1.16) has two singular points: 00 and $\infty$. The singularity at 00 is regular (Fuchsian), and the solution obtained by the well-known Frobenius method is the ${}_{0}F_{1}$ hypergeometric function, which we denote $F_{\alpha}$. We usually prefer its Olver normalized version ${\bf F}_{\alpha}:=\frac{F_{\alpha}}{\Gamma(\alpha+1)}$, closely related to the Bessel function, both standard and modified.
Another standard solution of the ${}_{0}F_{1}$ equation, corresponding to the irregular singularity at $\infty$, is the function that we denote $U_{\alpha}$, This function is perhaps less known. Up to a coefficient, it coincides with the Meijer $G$-function $G_{0,2}^{2,0}(-;0,-\alpha;z)$. The function $U_{\alpha}$ is closely related to the Macdonald and Hankel functions.
In our paper, we treat ${\bf F}_{\alpha}$ and $U_{\alpha}$ functions as basic elements of our description of Bessel potentials. In our opinion, they are much more convenient for this purpose, rather than functions from the Bessel family, as it is done in the conventional treatment of this topic. The corresponding formulas are simpler and more transparent. This is especially visible when we consider non-Euclidean signatures, where the formulas involve analytic continuation across two branches and an irregular distribution at the junction of these branches. The ${\bf F}_{\alpha}$ and $U_{\alpha}$ functions are also convenient to see the transition from the Minkowski space to the deSitter and the universal cover of the AntideSitter space, as discussed in [13]. In fact, on the Minkowski space retarded/advanced and Feynman/anti-Feynman Bessel potentials are expressed in terms of ${\bf F}_{\alpha}$ and $U_{\alpha}$, and on the deSitter and Anti-deSitter space we need closely related Gegenbauer functions instead.
We also believe that there are some important novel features in our presentation of the Lorentzian case, which is tailored to the needs of Quantum Field Theory. In our opinion, it is quite remarkable how rich is the theory of Bessel potentials in the Lorentzian signature. We have four distinct Lorentz invariant Green functions of the Klein-Gordon equation, with important applications in physics. If we include also a few useful distinguished solutions to the Klein-Gordon equation (such as the Pauli-Jordan propagator, positive and negative frequency solution), then we obtain a whole menagerie of functions.
In our discussion we cover not only the massive and massless case, but also the tachyonic case. This case is quite curious, even though usually ignored in the physics literature. We also discuss identity (1.14), true for $m^{2}\geq 0$, but wrong in the tachyonic case. Remarkably, this identity sometimes, but not always, generalizes to curved spacetimes, as analyzed recently in [13].
In our treatment, we pay special attention to the distributional character of Bessel potentials. This is unproblematic in the Euclidean signature, where Bessel potentials are given by (locally) integrable functions. This is not the case in non-Euclidean signatures. In particular, it is interesting to look at the functions ${\bf F}_{\alpha}$ and $U_{\alpha}$ as defining distributions on the real line. With this interpretation in mind, well-known identities have to be reformulated, see e.g. (2.47).
Finally, let us mention that there exist a large literature about Green functions of the Klein-Gordon equation on curved spacetimes. In the generic context their explicit expression is not possible, and often instead of exact Green functions one restricts oneself to parametrices, that is inverses modulo smoothing terms. The existence of exactly four parametrices that generalize $G^{\mathrm{F}/\overline{\mathrm{F}}}$ and $G^{\vee/\wedge}$ is the result of a famous paper by Duistermaat and Hörmander [14]. It is also remarkable that expansions similar to (5.34)-(5.37) describe singular parts of these parametrices also in curved spacetimes, where they can be derived from the Hadamard recursion relations (see Chapter 4 of [15] or Chapter 2 of [16].) The universality of these singular parts is an important idea in Quantum Field Theory on curved spacetimes [7].
2 Special functions related to the ${}_{0}F_{1}$ equation
2.1 The ${}_{0}F_{1}$ equation
Our presentation of Bessel potentials will use extensively ${}_{0}F_{1}$ hypergeometric functions, closely related to functions from the Bessel family. Surprisingly, they are seldom used and discussed in the literature. Therefore, we devote this section to a concise exposition of their properties, mostly following [17] and [18]. In particular, we will treat these functions as distributions on the real line, as explained in section 2.5, which leads to useful distributional identities which we have not seen in the literature.
Let $c\in{\mathbb{C}}$. The ${}_{0}F_{1}$ equation is
$(z{\rm\partial}_{z}^{2}+c{\rm\partial}_{z}-1)v(z)=0.$ (2.1)
If $c\neq 0,-1,-2,\dots$, then the only solution of the ${}_{0}F_{1}$ equation equal to $1$ at $z=0$ is called the ${}_{0}F_{1}$ hypergeometric function:
$F(c;z):=\sum_{j=0}^{\infty}\frac{1}{(c)_{j}}\frac{z^{j}}{j!},$
where $(c)_{j}$ denotes the Pochhammer symbol:
$(a)_{0}=1,$
$(a)_{n}:=a(a+1)\dots(a+n-1),$ $n=1,2,\dots$
$(a)_{n}:=\frac{1}{(a-n)\dots(a-1)},$ $n=\dots,-2,-1.$
$F(c;z)$ is defined for $c\neq 0,-1,-2,\dots$. Sometimes it is more convenient to consider the function
${\bf F}(c;z):=\frac{F(c;z)}{\Gamma(c)}=\sum_{j=0}^{\infty}\frac{1}{\Gamma(c+j)}\frac{z^{j}}{j!}$
defined for all $c$. For all parameters, we have an integral representation called the Schläfli formula:
$\frac{1}{2\pi{\rm i}}\int\limits_{]-\infty,0^{+},-\infty[}{\rm e}^{t}{\rm e}^{\frac{z}{t}}t^{-c}{\rm d}t$ $=$ ${\bf F}(c,z),\ \ \ \ {\rm Re}z>0,$
where the contour $]-\infty,0^{+},-\infty[$ starts at $\infty$- ∞, goes around 00 counterclockwise and returns to $\infty$- ∞.
Instead of $c$ it is often more natural to use $\alpha:=c-1$. Thus, we denote
$F_{\alpha}(z):=F(\alpha+1;z),\quad{\bf F}_{\alpha}(z):={\bf F}(\alpha+1;z).$ (2.2)
The following function is also a solution of the ${}_{0}F_{1}$ equation (1.16):
$U_{\alpha}(z)$ ::=:= ${\rm e}^{-2\sqrt{z}}z^{-\frac{\alpha}{2}-\frac{1}{4}}{}_{2}F_{0}\Big{(}\frac{1}{2}+\alpha,\frac{1}{2}-\alpha;-;-\frac{1}{4\sqrt{z}}\Big{)},$
where we used the ${}_{2}F_{0}$ function, see e.g. [17, 18]. $U_{\alpha}$ is a multivalued function. When talking about multivalued functions, we will usually consider their principal branches on the domain \${\mathbb{C}}\backslash]-\infty,0]$.
The function $U_{\alpha}$ rarely appears in the literature, except as a special case of Meijer’s function, see (2.32) below. Typically, it is represented through Macdonald or Hankel functions, which we describe further in equations (2.35), (2.37), and (2.36). In our opinion, however, the function $U_{\alpha}$ is often more convenient than Macdonald or Hankel functions.
$U_{\alpha}(z)$ has a symmetry
$U_{\alpha}(z)=z^{-\alpha}U_{-\alpha}(z).$ (2.3)
Alternatively, the function $U_{\alpha}$ can be defined by the integral representations valid for all $\alpha$:
$\frac{1}{\sqrt{\pi}}\int_{0}^{\infty}{\rm e}^{-t}{\rm e}^{-\frac{z}{t}}t^{-\alpha-1}{\rm d}t$ $=$ $U_{\alpha}(z),\ \ \ \ {\rm Re}z>0.$ (2.4)
For further reference, it is convenient to rewrite (2.4) as follows: For ${\rm Re}(m)>0$, we have
$\int_{0}^{\infty}{\rm e}^{-tm^{2}-\frac{x^{2}}{4t}}t^{-\alpha-1}{\rm d}t$ $=\sqrt{\pi}m^{2\alpha}U_{\alpha}\Big{(}\frac{m^{2}x^{2}}{4}\Big{)}.$ (2.5)
For ${\rm Re}(m)\geq 0$ (2.5) is still true in the sense of oscillatory integrals. By substituting $x^{2}\mapsto{\rm e}^{\pm{\rm i}\frac{\pi}{2}}x^{2},m^{2}\mapsto{\rm e}^{\pm{\rm i}\frac{\pi}{2}}m^{2},$ into (2.5) we obtain a pair of identities valid in terms of oscillatory integrals for $m>0$:
$\int_{0}^{\infty}{\rm e}^{\mp{\rm i}tm^{2}\mp\frac{x^{2}}{4t}}t^{-\alpha-1}{\rm d}t$ $={\rm e}^{{\rm i}\frac{\pi\alpha}{2}}\sqrt{\pi}m^{2\alpha}U_{\alpha}\Big{(}{\rm e}^{\pm{\rm i}\pi}\frac{m^{2}x^{2}}{4}\Big{)}.$ (2.6)
As $|z|\to\infty$ and $|\arg z|<2\pi-\epsilon$, $\epsilon>0$, we have
$U_{\alpha}(z)\sim{\rm exp}(-2z^{\frac{1}{2}})z^{-\frac{\alpha}{2}-\frac{1}{4}}.$ (2.7)
$U_{\alpha}$ is the unique solution of (1.16) with this property. (Note that the validity of (2.7) extends beyond $|\arg z|<\pi$, that is, beyond the principal sheet of the Riemann surface.)
We can express $U_{\alpha}$ in terms of the solutions of with a simple behavior at zero
$U_{\alpha}(z)$ $=$ $\frac{\sqrt{\pi}}{\sin\pi(-\alpha)}{\bf F}_{\alpha}(z)+\frac{\sqrt{\pi}}{\sin\pi\alpha}z^{-\alpha}{\bf F}_{-\alpha}(z).$ (2.8)
Alternatively, we can use the $U_{\alpha}$ function and its analytic continuation around 00 in the clockwise or anti-clockwise direction as the basis of solutions:
${\bf F}_{\alpha}(z)=\frac{\mp{\rm i}}{2\sqrt{\pi}}\left({\rm e}^{\mp{\rm i}\pi\alpha}U_{\alpha}(z)-{\rm e}^{\pm{\rm i}\pi\alpha}U_{\alpha}({\rm e}^{\pm{\rm i}2\pi}z)\right).$ (2.9)
Here is a version of (2.9) adapted to some applications:
${\bf F}_{\alpha}(-z)$ $=\frac{{\rm i}}{2\sqrt{\pi}}\left({\rm e}^{{\rm i}\pi\alpha}U_{\alpha}({\rm e}^{{\rm i}\pi}z)-{\rm e}^{-{\rm i}\pi\alpha}U_{\alpha}({\rm e}^{-{\rm i}\pi}z)\right),$ (2.10)
$z^{-\alpha}{\bf F}_{-\alpha}(-z)$ $=\frac{{\rm i}}{2\sqrt{\pi}}\left(U_{\alpha}({\rm e}^{{\rm i}\pi}z)-U_{\alpha}({\rm e}^{-{\rm i}\pi}z)\right).$ (2.11)
We have the recurrence relations
${\rm\partial}_{z}{\bf F}_{\alpha}(z)$ $={\bf F}_{\alpha+1}(z),$ (2.12)
$\left(z{\rm\partial}_{z}+\alpha\right){\bf F}_{\alpha}(z)$ $={\bf F}_{\alpha-1}(z);$ (2.13)
${\rm\partial}_{z}U_{\alpha}(z)$ $=-U_{\alpha+1}(z),$ (2.14)
$\left(z{\rm\partial}_{z}+\alpha\right)U_{\alpha}(z)$ $=-U_{\alpha-1}(z).$ (2.15)
$\alpha=m\in{\mathbb{Z}}$ is the degenerate case of the ${}_{0}F_{1}$ equation at 00. We have then
${\bf F}_{m}(z)=\sum_{n=\max(0,-m)}\frac{1}{n!(m+n)!}z^{n}.$
This easily implies the identity
${\bf F}_{m}(z)$ $=$ $z^{-m}{\bf F}_{-m}(z).$ (2.16)
In the degenerate case $U_{\alpha}(z)$ needs to be reexpressed using the de l’Hospital formula:
$U_{m}(z)$ $=$ $\frac{(-1)^{m+1}}{\sqrt{\pi}}\biggl{(}\sum_{k=1}^{m}\frac{(-1)^{k-1}(k-1)!}{(m-k)!}z^{-k}$
$+\sum_{j=0}^{\infty}\frac{\ln(z)-\psi(j+m+1)-\psi(j+1)}{j!(m+j)!}z^{j}\biggr{)}.$
In the degenerate case, the integral representation simplifies yielding the so-called Bessel integral representation. Besides, we have a generating function:
$\frac{1}{2\pi{\rm i}}\int\limits_{[0^{+}]}{\rm e}^{t+\frac{z}{t}}t^{-m-1}{\rm d}t$ $=$ ${\bf F}_{m}(z)=z^{-m}{\bf F}_{-m}(z),$
${\rm e}^{t}{\rm e}^{\frac{z}{t}}$ $=$ $\sum_{m\in{\mathbb{Z}}}t^{m}{\bf F}_{m}(z).$
Above, $[0^{+}]$ denotes the contour encircling 00 in the counterclockwise direction.
In the half-integer case, we can express the ${}_{0}F_{1}$ function in terms of elementary functions. Indeed,
$F_{-\frac{1}{2}}(z)=\cosh 2\sqrt{z},$ $U_{-\frac{1}{2}}(z)=\exp(-2\sqrt{z}),$ (2.18)
$F_{\frac{1}{2}}(z)=\frac{\sinh 2\sqrt{z}}{2\sqrt{z}},$ $U_{\frac{1}{2}}(z)=\frac{\exp(-2\sqrt{z})}{\sqrt{z}},$ (2.19)
and by the recurrence relations, we have for $k\in{\mathbb{N}}$
$F_{-\frac{1}{2}-k}(z)=$ $z^{k+\frac{1}{2}}\partial_{z}^{k}\Big{(}\frac{\cosh(2\sqrt{z}\big{)}}{\sqrt{z}}\Big{)},$ (2.20)
$F_{\frac{1}{2}+k}(z)=$ $\partial_{z}^{k}\Big{(}\frac{\sinh(2\sqrt{z})}{2\sqrt{z}}\Big{)},$ (2.21)
$U_{-\frac{1}{2}-k}(z)=$ $(-1)^{k}z^{k+\frac{1}{2}}\partial_{z}^{k}\Big{(}\frac{\exp(-2\sqrt{z})}{\sqrt{z}}\Big{)},$ (2.22)
$U_{\frac{1}{2}+k}(z)=$ $(-1)^{k}\partial_{z}^{k}\Big{(}\frac{\exp(-2\sqrt{z})}{\sqrt{z}}\Big{)}.$ (2.23)
2.2 Relationship to confluent functions
Recall that the confluent equation is
$(w\partial_{w}^{2}+(c-w)\partial_{w}-c)f(w)=0.$ (2.24)
Its standard solutions are
[…]
The ${}_{0}F_{1}$ equation can be reduced to a special class of the confluent equation by the so-called Kummer’s 2nd transformation:
$z{\rm\partial}_{z}^{2}+(\alpha+1){\rm\partial}_{z}-1$ (2.25)
$=$ $\frac{4}{w}{\rm e}^{-w/2}\Big{(}w\partial_{w}^{2}+(2\alpha+1-w)\partial_{w}-\alpha-\frac{1}{2}\Big{)}{\rm e}^{w/2},$ (2.26)
where $w=\pm 4\sqrt{z}$, $z=\frac{1}{16}w^{2}$. $F_{\alpha}$ and $U_{\alpha}$ can be expressed in terms of Kummer’s and Tricomi’s confluent function as follows:
$F_{\alpha}(z)$ $={\rm e}^{\mp 2\sqrt{z}}{}_{1}F_{1}\Big{(}\alpha+\frac{1}{2},2\alpha+1,\pm 4\sqrt{z}\Big{)},$ (2.27)
$U_{\alpha}(z)$ $=\frac{{\rm e}^{-2\sqrt{z}}}{2^{2\alpha+1}}U\Big{(}\alpha+\frac{1}{2},2\alpha+1,4\sqrt{z}\Big{)}.$ (2.28)
2.3 Relationship to Meijer G-functions
Solutions of hypergeometric equations ${}_{p}F_{q}$ can be expressed in terms of Meijer $G$-functions [19]. In particular, the ${}_{0}F_{1}$ equation can be solved by two distinguished functions
$G_{0,2}^{1,0}\Big{(}\begin{matrix}\\ 0,-\alpha\end{matrix}\Big{|}-z\Big{)}$ $:=\frac{1}{2\pi{\rm i}}\int_{L_{1}}\frac{\Gamma(-s){\rm e}^{{\rm i}\pi s}}{\Gamma(\alpha+1+s)}z^{s}{\rm d}s,$ (2.29)
$G_{0,2}^{2,0}\Big{(}\begin{matrix}\\ 0,-\alpha\end{matrix}\Big{|}z\Big{)}$ $:=\frac{1}{2\pi{\rm i}}\int_{L_{2}}\Gamma(-s)\Gamma(-\alpha-s)z^{s}{\rm d}s.$ (2.30)
Here, the contour $L_{1}$ goes from +$\infty$+ ∞ to +$\infty$+ ∞ and encircles ${\mathbb{N}}_{0}$, and the contour $L_{2}$ also goes from +$\infty$+ ∞ to +$\infty$+ ∞ and encircles ${\mathbb{N}}_{0}\cup({\mathbb{N}}_{0}-\alpha)$, both counterclockwise. Computing the residues and using the connection formula (2.8) we obtain
${\bf F}_{\alpha}(z)$ $=G_{0,2}^{1,0}\Big{(}\begin{matrix}\\ 0,-\alpha\end{matrix}\Big{|}-z\Big{)},$ (2.31)
$U_{\alpha}(z)$ $=\frac{1}{\sqrt{\pi}}G_{0,2}^{2,0}\Big{(}\begin{matrix}\\ 0,-\alpha\end{matrix}\Big{|}z\Big{)}.$ (2.32)
2.4 Relationship to Bessel functions
In the literature, the ${}_{0}F_{1}$ equation is seldom used. Much more frequent is the modified Bessel equation, which is equivalent to the ${}_{0}F_{1}$ equation. It is given by the operator
$z^{\frac{\alpha}{2}}\big{(}z{\rm\partial}_{z}^{2}+(\alpha+1){\rm\partial}_{z}-1\big{)}z^{-\frac{\alpha}{2}}$ $=$ $\partial_{w}^{2}+\frac{1}{w}\partial_{w}-1-\frac{\alpha^{2}}{w^{2}},$
where $z=\frac{w^{2}}{4}$, $w=\pm 2\sqrt{z}$.
Even more frequent is the (standard) Bessel equation given by:
$-z^{\frac{\alpha}{2}}\big{(}z{\rm\partial}_{z}^{2}+(\alpha+1){\rm\partial}_{z}-1\big{)}z^{-\frac{\alpha}{2}}$ $=$ $\partial_{u}^{2}+\frac{1}{u}\partial_{u}+1-\frac{\alpha^{2}}{u^{2}},$
where $z=-\frac{u^{2}}{4}$, $u=\pm 2{\rm i}\sqrt{z}$. Clearly, we can pass from the modified Bessel to the Bessel equation by $w=\pm{\rm i}u$.
The function ${\bf F}_{\alpha}$ is also seldom used. Instead, one uses the modified Bessel function and, even more frequently, the Bessel function:
$I_{\alpha}(w)$ $=$ $\Big{(}\frac{w}{2}\Big{)}^{\alpha}{\bf F}_{\alpha}\Big{(}\frac{w^{2}}{4}\Big{)},$ (2.33)
$J_{\alpha}(w)$ $=$ $\Big{(}\frac{w}{2}\Big{)}^{\alpha}{\bf F}_{\alpha}\Big{(}-\frac{w^{2}}{4}\Big{)}.$ (2.34)
They solve the modified Bessel, resp. the Bessel equation.
Instead of the $U_{\alpha}$ function one uses the Macdonald function, solving the modified Bessel equation:
$K_{\alpha}(w)$ $=$ $\frac{\sqrt{\pi}}{2}\Big{(}\frac{w}{2}\Big{)}^{\alpha}U_{\alpha}\Big{(}\frac{w^{2}}{4}\Big{)},$ (2.35)
and the Hankel functions of the 1st and 2nd kind, solving the Bessel equation:
$H_{\alpha}^{(1)}(w)=H_{\alpha}^{+}(w)$ $=$ $\frac{-{\rm i}}{\sqrt{\pi}}\Big{(}\frac{{\rm e}^{-{\rm i}\pi}w}{2}\Big{)}^{\alpha}U_{\alpha}\Big{(}{\rm e}^{-{\rm i}\pi}\frac{w^{2}}{4}\Big{)},$ (2.36)
$H_{\alpha}^{(2)}(w)=H_{\alpha}^{-}(w)$ $=$ $\frac{{\rm i}}{\sqrt{\pi}}\Big{(}\frac{{\rm e}^{{\rm i}\pi}w}{2}\Big{)}^{\alpha}U_{\alpha}\Big{(}{\rm e}^{{\rm i}\pi}\frac{w^{2}}{4}\Big{)}.$ (2.37)
Here are the relations between various functions from the Bessel family:
$H_{\alpha}^{\pm}(z)$ $=\frac{2}{\pi}{\rm e}^{\mp{\rm i}\frac{\pi}{2}(\alpha+1)}K_{\alpha}(\mp{\rm i}z),$ (2.38)
$H_{-\alpha}^{\pm}(z)$ $={\rm e}^{\pm\alpha\pi{\rm i}}H_{\alpha}^{\pm}(z),$ (2.39)
$J_{\alpha}(z)$ $=\frac{1}{2}\left(H_{\alpha}^{+}(z)+H_{\alpha}^{-}(z)\right),$ (2.40)
$I_{\alpha}(z)$ $=\frac{1}{\pi}\bigl{(}\mp{\rm i}K_{\alpha}({\rm e}^{\mp{\rm i}\pi}z)\pm{\rm i}{\rm e}^{{\rm i}\pi m}K_{\alpha}(z)\bigr{)}.$ (2.41)
2.5 ${\bf F}_{\alpha}$ and $U_{\alpha}$ functions as distributions
The function $U_{\alpha}(z)$ (and many others that we consider in this paper) are multivalued analytic functions defined on the Riemann surface of the logarithm. It has its principal branch on \${\mathbb{C}}\backslash]-\infty,0]$. For its analytic continuation around 00 we will often use the self-explanatory notation $U_{\alpha}({\rm e}^{{\rm i}\phi}z)$, where $z\in{\mathbb{C}}\backslash]-\infty,0]$ and $\phi\in{\mathbb{R}}$.
[…]
Thus we introduce the distribution on the real line
$U_{\alpha}({w}\pm{\rm i}0):=\lim_{\epsilon\searrow 0}U_{\alpha}({w}\pm{\rm i}\epsilon),$ (2.42)
where the right-hand side should be understood as the limit in the distributional sense. Note that for ${w}\neq 0$ these distributions are regular (in the sense of Appendix A.2) and given by analytic functions:
[…]
This easily follows from (2.8) and (2.1).
Recall that for $\alpha\not\in{\mathbb{N}}$ the symbol $w_{-}^{-\alpha}$ defined in (A.9) denotes the standard regularization of $|w|^{-\alpha}\theta(-w)$. The identity (2.11) for $\0w\in{\mathbb{R}}\backslash\{0\}$ can be rewritten as
[…]
(Note that both sides of (2.47) are zero for $w>0$). It is easy to see that for $\alpha\not\in{\mathbb{N}}$ (2.47) is a correct distributional identity, where the lhs is the product of the distribution $w_{-}^{-\alpha}$ and of the smooth function ${\bf F}_{-\alpha}(w)$, whereas the rhs is a linear combination of distributions defined in (2.42). (2.47) can be decomposed into a singular and regular part as follows:
$w_{-}^{-\alpha}{\bf F}_{-\alpha}(w)$ $=\sum_{j=0}^{\lfloor{\rm Re}\alpha\rfloor-1}\frac{w_{-}^{-\alpha+j}(-1)^{j}}{\Gamma(-\alpha+j+1)j!}+\sum_{j=\lfloor{\rm Re}\alpha\rfloor}^{\infty}\frac{w_{-}^{-\alpha+j}(-1)^{j}}{\Gamma(-\alpha+j+1)j!}$ (2.48)
The rhs of (2.47) is well-defined also for $\alpha\in{\mathbb{N}}$. We will define for such $\alpha$ the symbol on the lhs of (2.47) by the rhs. Using (A.14) for $\alpha\in{\mathbb{N}}$ we can thus write
$w_{-}^{-\alpha}{\bf F}_{-\alpha}(w)$ $=(-1)^{\alpha+1}\sum_{j=0}^{\alpha-1}\frac{(-1)^{j}\delta^{(\alpha-1-j)}(w)}{j!}+(-1)^{\alpha}{\bf F}_{\alpha}(w)\theta(-w).$ (2.49)
(Compare with (2.16), where you do not see the distributions supported at zero).
[…]
$K_{\alpha}\big{(}\sqrt{w\mp{\rm i}0}\big{)}$ :=$:=\begin{cases}K_{\alpha}\big{(}\sqrt{w}\big{)},&w>0,\\ K_{\alpha}(\mp{\rm i}\sqrt{-w}\big{)}=\pm{\rm i}\frac{\pi}{2}{\rm e}^{\pm{\rm i}\pi\alpha}H_{\alpha}^{\pm}\big{(}\sqrt{-w}\big{)},&w<0;\end{cases}$ (2.50a)
$H_{\alpha}^{\pm}\big{(}\sqrt{-w\pm{\rm i}0}\big{)}$ :=$:=\begin{cases}H_{\alpha}^{\pm}\big{(}\pm{\rm i}\sqrt{w}\big{)}=\mp{\rm i}\frac{2}{\pi}{\rm e}^{\mp{\rm i}\pi\alpha}K_{\alpha}\big{(}\sqrt{w}\big{)},&w>0,\\ H_{\alpha}^{\pm}\big{(}\sqrt{-w}\big{)},&w<0.\end{cases}$ (2.50b)
We believe, however, that it is more convenient in such situations to use the function $U_{\alpha}$. Indeed, we have
$U_{\alpha}\Big{(}\frac{w\mp{\rm i}0}{4}\Big{)}$ =$=\begin{cases}\frac{2^{\alpha+1}}{\sqrt{\pi}}(w\mp{\rm i}0)^{-\frac{\alpha}{2}}K_{\alpha}\big{(}\sqrt{w\mp{\rm i}0}\big{)}\\ \pm{\rm i}2^{\alpha}\sqrt{\pi}(w\mp{\rm i}0)^{-\frac{\alpha}{2}}H_{\alpha}^{\pm}\big{(}\sqrt{-w\pm{\rm i}0}\big{)}.\end{cases}$ (2.51)
3 Euclidean and anti-Euclidean signature
This section is devoted to Bessel potentials on the Euclidean space ${\mathbb{R}}^{d}$. $|x|:=\sqrt{x^{2}}$ will denote the Euclidean norm of $x\in{\mathbb{R}}^{d}$.
In this section, we will provide various expressions both in terms of the Bessel family functions $I_{\alpha},J_{\alpha},K_{\alpha},H_{\alpha}^{\pm}$, as well as in terms of the hypergeometric functions $F_{\alpha},U_{\alpha}$.
3.1 General exponents–Euclidean case
Consider first the Euclidean signature. For $m>0$ and ${\rm Re}\mu>0$ the function $\frac{1}{(p^{2}+m^{2})^{\frac{\mu}{2}}}$ defines a tempered distribution, hence one can compute its Fourier transform:
Theorem 1
Let $m>0$.
$G_{\mu,m}(x)$ $=\int\frac{{\rm e}^{{\rm i}px}}{(p^{2}+m^{2})^{\frac{\mu}{2}}}\frac{{\rm d}p}{(2\pi)^{d}}$ (3.1)
$=\frac{2}{\Gamma(\frac{\mu}{2})(4\pi)^{\frac{d}{2}}}\Big{(}\frac{|x|}{2m}\Big{)}^{\frac{\mu-d}{2}}K_{\frac{d-\mu}{2}}(m|x|)$ (3.2)
$=\frac{\sqrt{\pi}m^{d-\mu}}{\Gamma(\frac{\mu}{2})(4\pi)^{\frac{d}{2}}}U_{\frac{d-\mu}{2}}\Big{(}\frac{m^{2}x^{2}}{4}\Big{)}.$ (3.3)
Proof.  By (A.1),
$\frac{1}{(2\pi)^{d}}\int\frac{{\rm e}^{{\rm i}px}{\rm d}p}{(m^{2}+p^{2})^{\frac{\mu}{2}}}$ (3.4)
$=$ $\frac{1}{(2\pi)^{d}\Gamma(\frac{\mu}{2})}\int_{0}^{\infty}{\rm d}s\int{\rm d}ps^{\frac{\mu}{2}-1}{\rm e}^{-(m^{2}+p^{2})s}{\rm e}^{{\rm i}px}$ (3.5)
$=$ $\frac{1}{(4\pi)^{\frac{d}{2}}\Gamma(\frac{\mu}{2})}\int_{0}^{\infty}{\rm d}ss^{\frac{\mu}{2}-\frac{d}{2}-1}{\rm e}^{-m^{2}s-\frac{x^{2}}{4s}}$ (3.6)
Then we use (2.5). $\Box$
Note that the integrand of (3.1) is integrable for ${\rm Re}\mu>d$. Therefore, $G_{\mu,m}$ is bounded for such $\mu$. For instance,
$G_{\mu,m}(0)=\frac{1}{(2\pi)^{d}}\int\frac{{\rm d}p}{(p^{2}+m^{2})^{\frac{\mu}{2}}}$ $=\frac{m^{d-\mu}\Gamma(\frac{\mu-d}{2})}{(4\pi)^{\frac{d}{2}}\Gamma(\frac{\mu}{2})},\quad{\rm Re}\mu>d.$ (3.7)
3.2 General exponents–massless case
For $0<{\rm Re}\mu<d$ the following function is in $L_{\rm loc}^{1}({\mathbb{R}}^{d})$ and is bounded at infinity, hence it defines a regular distribution in ${\mathcal{S}}^{\prime}({\mathbb{R}}^{d})$:
$G_{\mu,0}(x)$ $:=\int\frac{{\rm e}^{{\rm i}px}}{|p|^{\mu}}\frac{{\rm d}p}{(2\pi)^{d}}$ (3.8)
$=\frac{\Gamma(\frac{d-\mu}{2})}{\Gamma(\frac{\mu}{2})(4\pi)^{\frac{d}{2}}}\Big{(}\frac{|x|}{2}\Big{)}^{\mu-d}.$ (3.9)
It is called the Riesz potential, and it is the massless limit of Bessel potentials:
Theorem 2
Let $0<{\rm Re}\mu<d$. Then
$G_{\mu,0}(x)=\lim_{m\to 0}G_{\mu,m}(x)$ (3.10)
in the sense of ${\mathcal{S}}^{\prime}({\mathbb{R}}^{d})$.
Proof.  One can prove this fact in the position space, see Subsection 4.2, where we give a proof in the case of a general signature. Instead, in this section we describe a proof based on the momentum space.
For $0<\mu<d$, $|p|^{-\mu}$ is a regular distribution. By using the Dominated Convergence Theorem we see that the pointwise limit
$\lim_{m\to 0}(p^{2}+m^{2})^{-\frac{\mu}{2}}=|p|^{-\mu}$ (3.11)
is a limit in the sense of ${\mathcal{S}}^{\prime}({\mathbb{R}}^{d})$. The Fourier transformation is a continuous operator on ${\mathcal{S}}^{\prime}({\mathbb{R}}^{d})$. Therefore, for considered $\mu$, (3.10) is true. $\Box$
3.3 General exponent–antiEuclidean case
Suppose now the scalar product is negative definite. For $m^{2}>0$, the function $\frac{1}{(-p^{2}+m^{2})^{\frac{\mu}{2}}}$ does not define uniquely a distribution, therefore one cannot compute its Fourier transform. However, if $m^{2}\in{\mathbb{C}}\backslash[0,\infty[$, then $\frac{1}{(-p^{2}+m^{2})^{\frac{\mu}{2}}}$ is a tempered distribution, and one can take its limit from above or below in the distributional sense:
$\frac{1}{(-p^{2}+m^{2}\pm{\rm i}0)^{\frac{\mu}{2}}}:=\lim_{\epsilon\searrow 0}\frac{1}{(-p^{2}+m^{2}\pm{\rm i}\epsilon)^{\frac{\mu}{2}}}.$ (3.12)
Thus we obtain two kinds of Bessel potentials in the antiEuclidean case:
Theorem 3
$G_{\mu,m}^{\mathrm{F}/\overline{\mathrm{F}}}(x)$ $=\int\frac{{\rm e}^{{\rm i}px}}{(-p^{2}+m^{2}\mp{\rm i}0)^{\frac{\mu}{2}}}\frac{{\rm d}p}{(2\pi)^{d}}$ (3.13)
$=\frac{\mp{\rm i}(\pm{\rm i})^{d}\pi}{\Gamma(\frac{\mu}{2})(4\pi)^{\frac{d}{2}}}\Big{(}\frac{|x|}{2m}\Big{)}^{\frac{\mu-d}{2}}H^{\mp}_{\frac{\mu-d}{2}}(m|x|)$ (3.14)
$=\frac{\mp{\rm i}{\rm e}^{\pm{\rm i}\frac{\pi\mu}{2}}\pi}{\Gamma(\frac{\mu}{2})(4\pi)^{\frac{d}{2}}}\Big{(}\frac{|x|}{2m}\Big{)}^{\frac{\mu-d}{2}}H^{\mp}_{\frac{d-\mu}{2}}(m|x|).$ (3.15)
$=\frac{{\rm e}^{\pm{\rm i}\pi\frac{d}{2}}\sqrt{\pi}m^{d-\mu}}{\Gamma(\frac{\mu}{2})(4\pi)^{\frac{d}{2}}}U_{\frac{d-\mu}{2}}\Big{(}\frac{{\rm e}^{\pm{\rm i}\pi}m^{2}x^{2}}{4}\Big{)}.$ (3.16)
Proof.  Using (A.2) and then (2.5) we obtain (3.16). $\Box$
Note that the Euclidean Bessel potential $G_{\mu,m}$ is well defined not only for $m\geq 0$, but also for ${\rm Re}(m)>0$, which guarantees $m^{2}\in{\mathbb{C}}\backslash]-\infty,0]$. Taking the limit at the imaginary line we can express the antiEuclidean Bessel potential in terms of the Euclidean one:
$G_{\mu,m}^{\mathrm{F}/\overline{\mathrm{F}}}(x)={\rm e}^{\mp{\rm i}\pi\frac{\mu}{2}}G_{\mu,\pm{\rm i}m}(x).$ (3.17)
3.4 Green functions of the Helmholtz equation
Bessel potentials with $\mu=2$ are Green functions of the Helmholtz equation
$(-E-\Delta)f(x)=g(x).$ (3.18)
More precisely, the Green function for $-E=m^{2}$ is
$G_{m}(x)$ $:=\int\frac{{\rm e}^{{\rm i}px}}{(p^{2}+m^{2})}\frac{{\rm d}p}{(2\pi)^{d}}$ (3.19)
$=\frac{1}{(2\pi)^{\frac{d}{2}}}\Big{(}\frac{|x|}{m}\Big{)}^{1-\frac{d}{2}}K_{\frac{d}{2}-1}(m|x|)$ (3.20)
$=\frac{\sqrt{\pi}m^{d-2}}{(4\pi)^{\frac{d}{2}}}U_{\frac{d}{2}-1}\Big{(}\frac{m^{2}x^{2}}{4}\Big{)}.$ (3.21)
and for $-E=-m^{2}$ we have two distinguished Green functions:
$G_{\mp{\rm i}m}(x)$ $=\int\frac{{\rm e}^{{\rm i}xp}}{(p^{2}-m^{2}\mp{\rm i}0)}\frac{{\rm d}p}{(2\pi)^{d}}$ (3.22)
=±$=\pm\frac{{\rm i}}{4}\Big{(}\frac{m}{2\pi|x|}\Big{)}^{\frac{d}{2}-1}H_{\frac{d}{2}-1}^{\pm}(m|x|)$ (3.23)
$=-(-{\rm i})^{d}\frac{\sqrt{\pi}m^{d-2}}{(4\pi)^{\frac{d}{2}}}U_{\frac{d}{2}-1}\Big{(}-\frac{m^{2}(x^{2}\pm{\rm i}0)}{4}\Big{)}.$ (3.24)
$G_{\mp{\rm i}m}(x)$ coincide with the case $\mu=2$ of the anti-Euclidean Bessel potential (3.13) multiplied by $-1$.
3.5 Averages of plane waves on sphere
Consider the sphere in ${\mathbb{R}}^{d}$ of radius $m$, denoted ${\mathbb{S}}_{m}^{d-1}={\mathbb{S}}_{m}$. Let ${\rm d}\Omega_{m}$ be the natural measure on ${\mathbb{S}}_{m}$. As an application of Bessel potentials, we will compute the Fourier transform of the measure on ${\mathbb{S}}_{m}$.
Theorem 4
$\int_{{\mathbb{S}}_{m}}{\rm e}^{{\rm i}px}{\rm d}\Omega_{m}(p)=$ $2m^{d-1}\pi^{\frac{d}{2}}{\bf F}_{\frac{d}{2}-1}\Big{(}-\frac{m^{2}x^{2}}{4}\Big{)}$ (3.25)
$=$ $m^{d-1}(2\pi)^{\frac{d}{2}}(m|x|)^{1-\frac{d}{2}}J_{\frac{d}{2}-1}(m|x|).$ (3.26)
Proof.  By the Sochocki-Plemejl formula, we have
$\delta(|p|-m)=2m\delta(p^{2}-m^{2})=\frac{2m}{2\pi{\rm i}}\Big{(}\frac{1}{p^{2}-m^{2}-{\rm i}0}-\frac{1}{p^{2}-m^{2}+{\rm i}0}\Big{)}.$ (3.27)
Therefore,
$\int_{{\mathbb{S}}_{m}}{\rm e}^{{\rm i}px}\Omega_{m}(p)=$ $\int{\rm e}^{{\rm i}px}\delta(|p|-m){\rm d}p$ (3.28)
$=$ $\frac{2m}{2\pi{\rm i}}\int{\rm e}^{{\rm i}px}\Big{(}\frac{1}{p^{2}-m^{2}-{\rm i}0}-\frac{1}{p^{2}-m^{2}+{\rm i}0}\Big{)}{\rm d}p$ (3.29)
$=$ $\frac{m(2\pi)^{d}}{\pi{\rm i}}\Big{(}G_{-{\rm i}m}(x)-G_{{\rm i}m}(x)\Big{)}$ (3.30)
$=m^{d-1}\pi^{\frac{d-1}{2}}$ $\Big{(}(-{\rm i})^{d-1}U_{\frac{d}{2}-1}\Big{(}\frac{{\rm e}^{-{\rm i}\pi}m^{2}x^{2}}{4}\Big{)}-{\rm i}^{d-1}U_{\frac{d}{2}-1}\Big{(}\frac{{\rm e}^{{\rm i}\pi}m^{2}x^{2}}{4}\Big{)}\Big{)}$ (3.31)
$=$ $2m^{d-1}\pi^{\frac{d}{2}}{\bf F}_{\frac{d}{2}-1}\Big{(}-\frac{m^{2}x^{2}}{4}\Big{)},$ (3.32)
where at the end we used (2.10). $\Box$
Consider a radial function ${\mathbb{R}}^{d}\ni p\mapsto f(|p|)$. Its Fourier transform is also radial. (3.25) yields the identity
$\int f(|p|){\rm e}^{-{\rm i}px}{\rm d}p=$ $2\pi^{\frac{d}{2}}\int_{0}^{\infty}f(k){\bf F}_{\frac{d}{2}-1}\Big{(}-\frac{k^{2}x^{2}}{4}\Big{)}k^{d-1}{\rm d}k$ (3.33)
$=$ $(2\pi)^{\frac{d}{2}}\int_{0}^{\infty}f(k)J_{\frac{d}{2}-1}(k|x|)(k|x|)^{-\frac{d}{2}+1}k^{d-1}{\rm d}k,$ (3.34)
where $k=|p|$ has the meaning of the length of $p$.
Using ${\bf F}_{-\frac{1}{2}}(-z)=\frac{\cos 2\sqrt{z}}{\sqrt{\pi}}$ and ${\bf F}_{\frac{1}{2}}(-z)=\frac{\sin 2\sqrt{z}}{\sqrt{\pi z}}$ we obtain the low dimensional cases of (3.33):
$\int f(|p|){\rm e}^{-{\rm i}px}{\rm d}p$ $=2\int_{0}^{\infty}f(k)\cos(k|x|){\rm d}k,\ \ \ \ d=1;$ (3.35)
$=2\pi\int_{0}^{\infty}f(k){\bf F}_{0}\Big{(}-\frac{k^{2}x^{2}}{4}\Big{)}{\rm d}k=2\pi\int_{0}^{\infty}f(k)kJ_{0}(k|x|){\rm d}k,\ \ \ \ d=2;$ (3.36)
$=4\pi\int_{0}^{\infty}f(k)k^{2}\frac{\sin(k|x|)}{k|x|}{\rm d}k,\ \ \ \ d=3.$ (3.37)
3.6 Integral representations of the $U_{\alpha}$ function
As an illustration of the usefulness of (3.33), we will derive a certain integral represention of $U_{\alpha}$.
Applying (3.33) to (3.3) we obtain
$2\int_{0}^{\infty}\frac{k^{d-1}{\rm d}k}{(k^{2}+1)^{\frac{\mu}{2}}}{\bf F}_{\frac{d}{2}-1}\Big{(}-\frac{r^{2}k^{2}}{4}\Big{)}=\frac{\sqrt{\pi}}{\Gamma(\frac{\mu}{2})}U_{\frac{d-\mu}{2}}\Big{(}\frac{r^{2}}{4}\Big{)}.$ (3.38)
Specifying $d=1$ and $d=3$ we obtain
$2\int_{0}^{\infty}\frac{\cos(kr)}{(k^{2}+1)^{\frac{\mu}{2}}}{\rm d}k$ $=\frac{\sqrt{\pi}}{\Gamma(\frac{\mu}{2})}U_{\frac{1-\mu}{2}}\Big{(}\frac{r^{2}}{4}\Big{)},$ (3.39)
$4\int_{0}^{\infty}\frac{k\sin(kr)}{(k^{2}+1)^{\frac{\mu}{2}}r}{\rm d}k$ $=\frac{\sqrt{\pi}}{\Gamma(\frac{\mu}{2})}U_{\frac{3-\mu}{2}}\Big{(}\frac{r^{2}}{4}\Big{)}.$ (3.40)
(3.40) could be also deduced from (3.39) by differentiating wrt $r$ and using the recurrence relation (2.14). Setting $\alpha=\frac{\mu-1}{2}$ in (3.39), we obtain the Poisson representation of the $U_{\alpha}$ function:
$U_{\alpha}\Big{(}\frac{r^{2}}{4}\Big{)}$ $=\frac{\Gamma(\frac{1}{2}-\alpha)}{\sqrt{\pi}}\int_{-\infty}^{\infty}{\rm e}^{-{\rm i}kr}(k^{2}+1)^{\alpha-\frac{1}{2}}{\rm d}k,\ \ \ \alpha<0.$ (3.41)
4 General signature
4.1 Positive mass
Consider now a pseudo-Euclidean space of general signature ${\mathbb{R}}^{q,d-q}$. $\frac{1}{(p^{2}+m^{2})^{\frac{\mu}{2}}}$ no longer defines a tempered distribution in the general signature. Just as in the antiEuclidean case, there are two natural regularizations of this function:
$\frac{1}{(p^{2}+m^{2}\pm{\rm i}0)^{\frac{\mu}{2}}}:=\lim_{\epsilon\searrow 0}\frac{1}{(p^{2}+m^{2}\pm{\rm i}\epsilon)^{\frac{\mu}{2}}}.$ (4.1)
They lead to two kinds of the Bessel potential:
Theorem 5
Let $m>0$ (or more generally ${\rm Re}(m)>0$). Then
$G_{\mu,m}^{\mathrm{F}/\overline{\mathrm{F}}}(x)=$ $\int\frac{{\rm e}^{{\rm i}px}}{(m^{2}+p^{2}\mp{\rm i}0)^{\frac{\mu}{2}}}\frac{{\rm d}p}{(2\pi)^{d}}$ (4.2)
$=$ $\frac{2(\pm{\rm i})^{q}}{\Gamma(\frac{\mu}{2})(4\pi)^{\frac{d}{2}}}\Big{(}\frac{\sqrt{x^{2}\pm{\rm i}0}}{2m}\Big{)}^{\frac{\mu-d}{2}}K_{\frac{d-\mu}{2}}\big{(}\sqrt{m^{2}(x^{2}\pm{\rm i}0)}\big{)}$ (4.3)
$=$ $\mp\frac{\pi{\rm i}(\pm{\rm i})^{q}}{\Gamma(\frac{\mu}{2})(4\pi)^{\frac{d}{2}}}\Big{(}\frac{\sqrt{x^{2}\pm{\rm i}0}}{2m}\Big{)}^{\frac{\mu-d}{2}}H_{\frac{\mu-d}{2}}^{\mp}\big{(}\sqrt{m^{2}({-x^{2}\mp{\rm i}0})}\big{)}$ (4.4)
$=$ (±$\frac{(\pm{\rm i})^{q}\sqrt{\pi}m^{d-\mu}}{\Gamma(\frac{\mu}{2})(4\pi)^{\frac{d}{2}}}U_{\frac{d-\mu}{2}}\Big{(}\frac{m^{2}(x^{2}\pm{\rm i}0)}{4}\Big{)}.$ (4.5)
Remark 6
[…]
Anyway, in our opinion the expression in terms of $U_{\alpha}$, (4.5), is preferable.
Proof of Thm 5. Using (A.2) and (A.4) we obtain
$\frac{1}{(2\pi)^{d}}\int\frac{{\rm e}^{{\rm i}px}{\rm d}p}{(m^{2}+p^{2}\mp{\rm i}0)^{\frac{\mu}{2}}}$
$=$ $\frac{{\rm e}^{\pm{\rm i}\frac{\pi\mu}{4}}}{(2\pi)^{d}\Gamma(\frac{\mu}{2})}\int_{0}^{\infty}{\rm d}t\int{\rm d}p{\rm e}^{\mp{\rm i}t(m^{2}+p^{2})}t^{\frac{\mu}{2}-1}{\rm e}^{{\rm i}px}$
$=$ (±$\frac{(\pm{\rm i})^{q}{\rm e}^{\pm{\rm i}\frac{\pi}{2}(\frac{\mu-d}{2})}\pi^{\frac{d}{2}}}{(4\pi)^{\frac{d}{2}}\Gamma(\frac{\mu}{2})}\int_{0}^{\infty}{\rm d}t{\rm e}^{\mp{\rm i}(tm^{2}-\frac{x^{2}}{4t})}t^{\frac{\mu-d}{2}-1}.$ (4.6)
Then we apply (2.6). $\Box$
4.2 Zero mass
For $0<{\rm Re}\mu<d$ let us introduce two distributions in ${\mathcal{S}}^{\prime}$
$G_{\mu,0}^{\mathrm{F}/\overline{\mathrm{F}}}(x)$ $:=\int\frac{{\rm e}^{{\rm i}px}}{(p^{2}\mp{\rm i}0)^{\frac{\mu}{2}}}\frac{{\rm d}p}{(2\pi)^{d}}$ (4.7)
=(±$=\frac{(\pm{\rm i})^{q}\Gamma(\frac{d-\mu}{2})}{\Gamma(\frac{\mu}{2})(4\pi)^{\frac{d}{2}}}\Big{(}\frac{x^{2}\pm{\rm i}0}{4}\Big{)}^{\frac{\mu-d}{2}}.$ (4.8)
They will be called Feynman/antiFeynman Riesz potentials. They are massless limits of the corresponding Bessel potentials:
Theorem 7
For $0<{\rm Re}\mu<d$ we have
$G_{\mu,0}^{\mathrm{F}/\overline{\mathrm{F}}}(x)=\lim_{m\searrow 0}G_{\mu,m}^{\mathrm{F}/\overline{\mathrm{F}}}(x)$ (4.9)
in the sense of ${\mathcal{S}}^{\prime}$.
Proof.  Surprisingly, a momentum space proof, from the Euclidean case, seems to be difficult to generalize to the non-Euclidean case. Instead, we will present a proof in the position space.
Using the decomposition (2.45) of the function $U_{\alpha}$, we can write
$G_{\mu,m}(x)$ =(±$=\frac{(\pm{\rm i})^{q}}{\Gamma(\frac{\mu}{2})(4\pi)^{\frac{d}{2}}}\Bigg{(}\sum_{j=0}^{\lfloor{\rm Re}\frac{d-\mu}{2}\rfloor-1}\frac{(-1)^{j}m^{j}\Gamma(\frac{d-\mu}{2}-j)}{j!}\Big{(}\frac{(x^{2}\pm{\rm i}0)}{4}\Big{)}^{j-\frac{d-\mu}{2}}$ (4.10)
$+m^{d-\mu}U_{\frac{d-\mu}{2}}^{\rm reg}\Big{(}\frac{m^{2}(x^{2}\pm{\rm i}0)}{4}\Big{)}\Bigg{)}.$ (4.11)
The line (4.10) obviously converges to (4.8). By (2.7), $U_{\frac{d-\mu}{2}}^{\rm reg}$ is a continuous function of a polynomial growth at infinity. Therefore, the second line (4.11) converges to zero in ${\mathcal{S}}^{\prime}$. $\Box$
Note that as a consequence of the above theorem and of the continuity of the Fourier transformation on ${\mathcal{S}}^{\prime}({\mathbb{R}}^{d})$ we can infer that
$\lim_{m\searrow 0}\frac{1}{(p^{2}+m^{2}\mp{\rm i}0)^{\frac{\mu}{2}}}=\frac{1}{(p^{2}\mp{\rm i}0)^{\frac{\mu}{2}}}$ (4.12)
in the sense of ${\mathcal{S}}^{\prime}$.
4.3 Scaling degree of distributions
Let us start by defining the action of a dilation by $\lambda$ on a distribution $T(x)$ as $T_{\lambda}(x)=T(\lambda x)$, by which we mean the dual action to the dilation on test functions
$\langle T_{\lambda}|f\rangle=\int T(\lambda x)f(x){\rm d}x=\lambda^{-d}\int T(x)f(\lambda^{-1}x){\rm d}x.$ (4.13)
Given a distribution $T\in\mathcal{D}^{\prime}({\mathbb{R}}^{d})$, we define its scaling degree $\mathrm{sd}(T)$ as
$\mathrm{sd}(T)=\inf\Big{\{}\omega:\lim_{\lambda\searrow 0}\lambda^{\omega}T_{\lambda}=0\ \text{ in }\mathcal{D}^{\prime}({\mathbb{R}}^{d})\Big{\}}.$ (4.14)
The scaling degree of a distribution is often used in mathematical analysis of Quantum Field Theory [6, 7].
Let us compute the scaling degree of Bessel potentials.
Theorem 8
$\mathrm{sd}G_{m,\mu}^{\mathrm{F}/\overline{\mathrm{F}}}=\begin{cases}d-\mu,&0<\mu\leq d;\\ 0,&d\leq\mu.\end{cases}.$ (4.15)
Proof.  For $0<\mu<d$, the Riesz potentials $G_{\mu,0}^{\mathrm{F}/\overline{\mathrm{F}}}$ defined in (4.8) are homogeneous:
$G_{\mu,0}^{\mathrm{F}/\overline{\mathrm{F}}}(\lambda x)=\lambda^{\mu-d}G_{\mu,0}^{\mathrm{F}/\overline{\mathrm{F}}}(x).$ (4.16)
So $\mathrm{sd}G_{\mu,0}^{\mathrm{F}/\overline{\mathrm{F}}}=d-\mu$.
By the definition of the Bessel potential, the mass dependence is (1.2)
$G_{\mu,m}(\lambda x)=\lambda^{\mu-d}G_{\mu,\lambda m}(x),$ (4.17)
so, according to Theorem 7,
$\lim_{\lambda\searrow 0}\lambda^{d-\mu}G_{\mu,m}(\lambda x)=\lim_{\lambda\searrow 0}G_{\mu,\lambda m}(x)=G_{\mu,0}(x),$ (4.18)
which shows that $\mathrm{sd}G_{\mu,m}^{\mathrm{F}/\overline{\mathrm{F}}}=d-\mu$ for any mass $m$ and $0<\mu<d$.
For $d<\mu$, $G_{m}^{F/\overline{\mathrm{F}}}$ is a continuous bounded function, so its scaling degree is 00.
For $d=\mu$, we have
$G_{d,m}(x)=\frac{(\pm{\rm i})^{q}\sqrt{\pi}m^{d-\mu}}{\Gamma(\frac{d}{2})(4\pi)^{\frac{d}{2}}}U_{0}\Big{(}\frac{m^{2}(x^{2}\pm{\rm i}0)}{4}\Big{)}.$ (4.19)
Now, we can use the bound (2.7) and the expansion (2.1)
[…]
where $H$ is an entire function, just as $\mathbf{F}_{0}$. Using this we easily show that for $\omega>0$
$\lambda^{\omega}G_{d,m}(\lambda x)\to 0$ (4.22)
in the sense of ${\mathcal{S}}^{\prime}$. $\Box$
5 The Minkowski signature
The Lorentzian signature is especially important, both because of its physical relevance and rich mathematical properties. The spaces ${\mathbb{R}}^{1,d-1}$ and ${\mathbb{R}}^{d-1,1}$ are two kinds of a Minkowski space, that is, a pseudo-Euclidean space with a Lorentzian signature. We will treat ${\mathbb{R}}^{1,d-1}$ as the standard form of a Minkowski space. $x^{0}$ will denote the first coordinate of ${\mathbb{R}}^{1,d-1}$, which we assume to be timelike (having a negative coefficient in the scalar product). The remaining, spacelike coordinates will be denoted $\vec{x}$, so that $x=(x^{0},\vec{x})$. In other words,
$x^{2}=-(x^{0})^{2}+\vec{x}^{2}=-(x^{0})^{2}+(x^{1})^{2}+\cdots+(x^{d-1})^{2}.$ (5.1)
The future and the past light cone will be denoted
$J^{\vee}$ :=$:=\{x\in{\mathbb{R}}^{1,d-1}\ :\quad x^{2}\leq 0,\quad x^{0}\geq 0\},$
$J^{\wedge}$ :=$:=\{x\in{\mathbb{R}}^{1,d-1}\ :\quad x^{2}\leq 0,\quad x^{0}\leq 0\}.$
In this section, we will only use the hypergeometric functions ${\bf F}_{\alpha},U_{\alpha}$.
5.1 General exponent
Let $m>0$. The set $m^{2}+p^{2}$ consists of two connected components: the future and the past mass hyperboloid. Therefore, the following four regularizations of $\frac{1}{(m^{2}+p^{2})^{\frac{\mu}{2}}}$ are tempered distributions invariant wrt the orthochronous Lorentz group:
$\frac{1}{(m^{2}+p^{2}\pm{\rm i}0)^{\frac{\mu}{2}}},\quad\frac{1}{(m^{2}+p^{2}\pm{\rm i}0{\rm sgn}p^{0})^{\frac{\mu}{2}}}.$ (5.2)
Their inverse Fourier transforms define four kinds of Bessel potentials:
$G_{\mu,m}^{\mathrm{F}/\overline{\mathrm{F}}}(x):=$ $\int\frac{{\rm e}^{{\rm i}px}}{(m^{2}+p^{2}\mp{\rm i}0)^{\frac{\mu}{2}}}\frac{{\rm d}p}{(2\pi)^{d}}$ (5.3)
$G_{\mu,m}^{\vee/\wedge}(x):=$ $\int\frac{{\rm e}^{{\rm i}px}}{(m^{2}+p^{2}\mp{\rm i}0{\rm sgn}p^{0})^{\frac{\mu}{2}}}\frac{{\rm d}p}{(2\pi)^{d}}.$ (5.4)
By the following well-known argument, found e.g. in various standard textbooks on quantum field theory, we can show that $G_{\mu,m}^{\vee/\wedge}$ have causal supports.
Theorem 9
${\rm supp}G_{\mu,m}^{\vee/\wedge}\subset J^{\vee/\wedge}$.
Proof.  For definiteness, consider (5.4) with the minus sign. In order to prove that its support is contained in $J^{\vee}$, by the Lorentz invariance it suffices to prove that it is zero for $x^{0}<0$. We write
$\int\frac{{\rm e}^{{\rm i}px}{\rm d}p}{(p^{2}+m^{2}-{\rm i}0{\rm sgn}p^{0})^{\frac{\mu}{2}}}=\int\frac{{\rm e}^{-{\rm i}p^{0}x^{0}+{\rm i}\vec{p}\vec{x}}{\rm d}p^{0}{\rm d}\vec{p}}{\big{(}\vec{p}^{2}+m^{2}-(p^{0}+{\rm i}0)^{2}\big{)}^{\frac{\mu}{2}}}.$
[…]
Theorem 10
We have the identity
$G_{\mu,m}^{\mathrm{F}}(x)+G_{\mu,m}^{\overline{\mathrm{F}}}(x)$ $=G_{\mu,m}^{\vee}(x)+G_{\mu,m}^{\wedge}(x)$ (5.5)
Here are the expressions for the Bessel potentials in the position space:
$G_{\mu,m}^{\mathrm{F}/\overline{\mathrm{F}}}(x)=$$\frac{\pm{\rm i}\sqrt{\pi}m^{d-\mu}}{\Gamma(\frac{\mu}{2})(4\pi)^{\frac{d}{2}}}U_{\frac{d-\mu}{2}}\Big{(}\frac{m^{2}(x^{2}\pm{\rm i}0)}{4}\Big{)}.$ (5.6)
$G_{\mu,m}^{\vee/\wedge}(x)=$ $\theta(\pm x^{0})\frac{2\pi}{\Gamma(\frac{\mu}{2})(4\pi)^{\frac{d}{2}}}\big{(}\tfrac{x^{2}}{4}\big{)}_{-}^{\frac{\mu-d}{2}}{\bf F}_{\frac{\mu-d}{2}}\Big{(}\frac{m^{2}x^{2}}{4}\Big{)}.$ (5.7)
where in (5.7) we used the notation introduced in (2.47).
Formula (5.7) involves the multiplication of a distribution by a discontinuous function, which in general is not well defined. At the end of this subsection we provide explain how this formula can be correctly interpreted.
Proof.  The identity (5.5) follows immediately from the defining formulas, that is from (5.3) and (5.4).
(5.6) is a special case of (4.5). Using (5.6) and (5.7) we obtain a simple expression for the sum of two Bessel potentials:
$G_{\mu,m}^{\vee}(x)+G_{\mu,m}^{\wedge}(x)=$ $\frac{-{\rm i}\sqrt{\pi}m^{d-\mu}}{\Gamma(\frac{\mu}{2})(4\pi)^{\frac{d}{2}}}\Bigg{(}U_{\frac{d-\mu}{2}}\Big{(}\frac{m^{2}x^{2}-{\rm i}0}{4}\Big{)}-U_{\frac{d-\mu}{2}}\Big{(}\frac{m^{2}x^{2}+{\rm i}0}{4}\Big{)}\Bigg{)}$ (5.8)
$=$ $\frac{2\pi}{\Gamma(\frac{\mu}{2})(4\pi)^{\frac{d}{2}}}\big{(}\tfrac{x^{2}}{4}\big{)}_{-}^{\frac{\mu-d}{2}}{\bf F}_{\frac{\mu-d}{2}}\Big{(}\frac{m^{2}x^{2}}{4}\Big{)},$ (5.9)
where again we used the notation introduced in (2.47). (5.9) is clearly supported in $J^{\wedge}\cup J^{\vee}$. By Thm 9, we know that $G_{\mu,m}^{\vee/\wedge}$ are supported in $J^{\vee/\wedge}$. Thus to find expressions for $G_{\mu,m}^{\vee/\wedge}$ we need to “split the distribution” (5.9) into two terms, one supported in $J^{\vee}$ and the other in $J^{\wedge}$.
Using Proposition 11 to justify the multiplication of a distribution (5.9) by the (discontinuous) function $\theta(\pm{x}^{0})$, we can define
$\tilde{G}_{\mu,m}^{\vee/\wedge}(x)$ $=\theta(\pm x^{0})\frac{2\pi}{\Gamma(\frac{\mu}{2})(4\pi)^{\frac{d}{2}}}\big{(}\tfrac{x^{2}}{4}\big{)}_{-}^{\frac{\mu-d}{2}}{\bf F}_{\frac{\mu-d}{2}}\Big{(}\frac{m^{2}x^{2}}{4}\Big{)}.$ (5.10)
Clearly, $\tilde{G}_{\mu,m}^{\vee/\wedge}$ are supported in $J^{\vee/\wedge}$. Besides,
$G_{\mu,m}^{\vee}(x)+G_{\mu,m}^{\wedge}(x)=\tilde{G}_{\mu,m}^{\vee}(x)+\tilde{G}_{\mu,m}^{\wedge}(x).$ (5.11)
But $J^{\vee}\cap J^{\wedge}=\{0\}$. Therefore, $G_{\mu,m}^{\vee/\wedge}-\tilde{G}_{\mu,m}^{\vee/\wedge}$ is a distribution supported in $\{0\}$, that is, a linear combination of $\delta^{(\alpha)}(x)$
$B_{\mu,m}^{\vee/\wedge}:=G_{\mu,m}^{\vee/\wedge}-\tilde{G}_{\mu,m}^{\vee/\wedge}=\sum_{|\alpha|<n}c_{\alpha,m}^{\vee/\wedge}\delta^{(\alpha)}(x).$ (5.12)
(5.11) implies $B_{\mu,m}^{\vee}(x)=-B_{\mu,m}^{\wedge}(x)$. The symmetry in $x\mapsto-x,\vee/\wedge\mapsto\wedge/\vee$ of (5.4) and (5.7) allows us to write $G_{\mu,m}^{\vee}(x)=G_{\mu,m}^{\wedge}(-x)$, $\tilde{G}_{\mu,m}^{\vee}(x)=\tilde{G}_{\mu,m}^{\wedge}(-x)$, and therefore $B_{\mu,m}^{\vee/\wedge}(x)=B_{\mu,m}^{\wedge/\vee}(-x)=-B_{\mu,m}^{\vee/\wedge}(-x).$ Its action on a test function $\phi\in{\mathcal{S}}({\mathbb{R}}^{1,d-1})$ is
$\langle B_{\mu,m}^{\vee/\wedge},\phi\rangle=\sum_{|\alpha|<n}(-1)^{|\alpha|}c_{\alpha,m}^{\vee/\wedge}(\partial_{x}^{\alpha}\phi)(0)=\sum_{|\alpha|<n}c_{\alpha,m}^{\vee/\wedge}(\partial_{-x}^{\alpha}\phi)(0),$ (5.13)
so $c_{\alpha,m}^{\vee/\wedge}=0$ for even $|\alpha|$. $G_{\mu,m}^{\vee/\wedge}$ and $\tilde{G}_{\mu,m}^{\vee/\wedge}$ are invariant with respect to the action of the proper Lorentz group. The same must apply to their difference $B_{\mu,m}^{\vee/\wedge}$. Derivatives evaluated at 00 transform as vectors under the action of the Lorentz group. However, $\langle B_{\mu,m}^{\vee/\wedge},\phi\rangle$ is a sum of terms with only odd number of indices so it cannot be invariant under the action of the Lorentz group unless $B_{\mu,m}^{\vee/\wedge}=0$. We conclude that $G_{\mu,m}^{\vee/\wedge}=\tilde{G}_{\mu,m}^{\vee/\wedge}$. $\Box$
$G_{\mu,m}^{\mathrm{F}}$ will be called the Feynman Bessel potential and $G_{\mu,m}^{\overline{\mathrm{F}}}$ the anti-Feynman Bessel potential. These names are somewhat artificial in the context of a general $\mu$. Their justification comes from the case $\mu=2$, where these Bessel potentials coincide with the Feynman and anti-Feynman propagator known from Quantum Field Theory.
The distribution $G_{\mu,m}^{\vee}$ will be called the forward or retarded Bessel potential, and $G_{\mu,m}^{\wedge}$ the backward or advanced Bessel potential.
For $0<{\rm Re}\mu<d$ we also have the massless Riesz potentials:
$G_{\mu,0}^{\mathrm{F}/\overline{\mathrm{F}}}(x)=$$\frac{\pm{\rm i}\Gamma(\frac{d-\mu}{2})}{\Gamma(\frac{\mu}{2})(4\pi)^{\frac{d}{2}}}\Big{(}\frac{x^{2}\pm{\rm i}0}{4}\Big{)}^{\frac{\mu-d}{2}},$ (5.14)
$G_{\mu,0}^{\vee/\wedge}(x)=$ $\theta(\pm x^{0})\frac{2\pi}{\Gamma(\frac{\mu}{2})\Gamma(\frac{\mu-d+2}{2})(4\pi)^{\frac{d}{2}}}\big{(}\tfrac{x^{2}}{4}\big{)}_{-}^{\frac{\mu-d}{2}}.$ (5.15)
As we mentioned above, the formula (5.7) for the advanced and retarded Bessel potential involves a product of two distributions, and therefore it needs a justfication. We will explain two approaches how to interpret this formula.
The first approach is quite elementary. It uses the identification ${\mathbb{R}}^{1,d-1}\simeq{\mathbb{R}}\times{\mathbb{R}}^{d-1}$, with the first variable denoted $x^{0}$ or $t$. For the remaining variables $\vec{x}$ we will later use spherical coordinates $(r,\Omega)$ with $r=|\vec{x}|$. For $n,m\in{\mathbb{N}}_{0}$ and $\chi\in{\mathcal{S}}({\mathbb{R}}^{1,d-1})$ let us introduce the semi-norms, which involve only the variables $\vec{x}\in{\mathbb{R}}^{d-1}$:
$\|\chi(t,\cdot)\|_{n,m}=\sup_{\vec{x}\in\mathbb{R}^{d-1},|\alpha|=n,|\beta|=m}=|\vec{x}^{\alpha}(\partial_{\vec{x}}^{\beta}\chi)(t,\vec{x})|.$
Proposition 11
Let $\text{Re}\nu<d$. Then there exist $c_{k}$, $k=0,\dots,\lfloor\frac{{\rm Re}\nu}{2}\rfloor$, such that for any $\phi\in{\mathcal{S}}({\mathbb{R}}^{1,d-1})$
$\Big{|}\int\big{(}x^{2}\big{)}_{-}^{-\frac{\nu}{2}}\phi(x){\rm d}x\Big{|}\leq\sum_{k=0}^{\lfloor\frac{{\rm Re}\nu}{2}\rfloor}\int c_{k}|t|^{d-{\rm Re}\nu+k-1}\|\phi(t,\cdot)\|_{0,k}{\rm d}t,$ (5.16)
where the coefficients $|t|^{d-{\rm Re}\nu+k-1}$ are locally integrable and polynomially bounded at infinity. Therefore, if $f\in L^{\infty}({\mathbb{R}})$, then $f(x^{0})\big{(}x^{2}\big{)}_{-}^{-\frac{\nu}{2}}$ defines a tempered distribution on ${\mathbb{R}}^{d}$.
Proof.   Action of $\big{(}x^{2}\big{)}_{-}^{-\frac{\nu}{2}}$ on test a function $\phi\in\mathcal{S}({\mathbb{R}}^{d})$ is
$\int\big{(}x^{2}\big{)}_{-}^{-\frac{\nu}{2}}\phi(x){\rm d}x=\int_{-\infty}^{\infty}{\rm d}t\int_{0}^{|t|}{\rm d}r\int_{\mathbb{S}^{d-2}}{\rm d}\Omega(r^{2}-t^{2})_{-}^{-\frac{\nu}{2}}\phi(t,r,\Omega)r^{d-2}.$
For simplicity let us consider only $t>0$. We can expand $\phi$ around $r=t$
$\phi(t,r,\Omega)=\sum_{k=0}^{m}\frac{(r-t)^{k}}{k!}\phi^{(k)}(t,t,\Omega)+(r-t)^{m+1}\psi(t,r,\Omega).$
with $m=\lfloor\frac{\text{Re}\nu}{2}\rfloor-1$, where $\phi^{(k)}$ denote derivatives with respect to the $r$ variable. Note that $|\psi(t,r,\Omega)|\leq(m+1)!\|\phi(t,\cdot)\|_{0,m+1}$. Let
$a_{m+1}:=$ $\int_{0}^{\infty}{\rm d}t\int_{0}^{t}{\rm d}r\int_{\mathbb{S}^{d-2}}{\rm d}\Omega(t-r)^{\beta}(r+t)^{-\frac{\nu}{2}}\psi(t,r,\Omega)r^{d-2},$
with $\beta=\lfloor\frac{\text{Re}\nu}{2}\rfloor-\frac{\nu}{2},\quad-1<\text{Re}\beta\leq 0$, is the integral of the locally integrable function. We see that it is well defined and
$|a_{m+1}|\leq\int_{0}^{\infty}{\rm d}t\int_{0}^{t}{\rm d}r\int_{\mathbb{S}^{d-2}}{\rm d}\Omega(t-r)^{\text{Re}\beta}(t+r)^{-\text{Re}\frac{\nu}{2}}|\psi(t,r,\Omega)|r^{d-2}$
$\leq(m+1)!\int_{0}^{\infty}{\rm d}t\|\phi(t,\cdot)\|_{0,m+1}t^{\lfloor\frac{\text{Re}\nu}{2}\rfloor+d-\text{Re}\nu-1}\int_{0}^{1}{\rm d}r^{\prime}(1-r^{\prime})^{\beta}(1+r^{\prime})^{-\frac{\text{Re}\nu}{2}}r^{\prime d-2}\int_{\mathbb{S}^{d-2}}{\rm d}\Omega$
$=:C(d,\nu,m+1)\int_{0}^{\infty}{\rm d}tt^{\lfloor\frac{\nu}{2}\rfloor+d-\text{Re}\nu-1}\|\phi(t,\cdot)\|_{0,m+1}.$
Next, we look at each term of the expansion of $\phi(t,r,\Omega)$ in $k$
$a_{k}=$ $\int_{0}^{\infty}{\rm d}t\int_{0}^{t}{\rm d}r\int_{\mathbb{S}^{d-2}}{\rm d}\Omega(r^{2}-t^{2})_{-}^{-\frac{\nu}{2}}\frac{(r-t)^{k}}{k!}\phi^{(k)}(t,t,\Omega)r^{d-2}$
$=\frac{(-1)^{k}}{k!}$ $\int_{0}^{\infty}{\rm d}t\int_{0}^{t}{\rm d}r(r-t)^{-\frac{\nu}{2}+k}_{-}(t+r)^{-\frac{\nu}{2}}r^{d-2}\int_{\mathbb{S}^{d-2}}{\rm d}\Omega\phi^{(k)}(t,t,\Omega).$
Here, $(t-r)_{-}^{-\frac{\nu}{2}+k}$ is the (irregular) distribution, defined by (A.12). It yields a finite expression:
$\int_{0}^{t}{\rm d}r(r-t)_{-}^{-\frac{\nu}{2}+k}(t+r)^{-\frac{\nu}{2}}r^{d-2}$
$=t^{d-\nu+k-1}$ $\int_{0}^{1}{\rm d}r^{\prime}(r^{\prime}-1)_{-}^{-\frac{\nu}{2}+k}(1+r^{\prime})^{-\frac{\nu}{2}}r^{d-2}=:t^{d-\nu+k-1}\tilde{C}(d,\nu,k).$
Because $d-\text{Re}\nu+k-1\geq d-\text{Re}\nu-2>-1$, dependence on $t$ is locally integrable and bounded by a polynomial. For $k=0,1,\ldots,m+1$ we can write
$|a_{k}|\leq C(d,\nu,k)\int_{0}^{\infty}{\rm d}t\ t^{d-\text{Re}\nu+k-1}\|\phi(t,\cdot)\|_{0,k}.$
For fixed $d,\nu$, we have the inequality (5.16) showing that homogeneous distributions are tempered distribution. $\Box$
Now we have $d-\mu<d$, and therefore Proposition 11 shows that we can multiply the distribution $\big{(}\tfrac{x^{2}}{4}\big{)}_{-}^{\frac{\mu-d}{2}}$ by the discontinuous but bounded function $\theta(\pm x^{0})$. The resulting distribution is then multiplied by the smooth function ${\bf F}_{\frac{\mu-d}{2}}\Big{(}\frac{m^{2}x^{2}}{4}\Big{)}$, obtaining the right hand side of (5.7).
An alternative way to define the product in (5.7) is based on the concept of the wave front set [20]. Here are the wave front sets of the distributions contained in (5.7):
$\mathrm{WF}\left(\theta(t)\right)=$$\Big{\{}\left((0,\vec{x}),(\tau,0)\right)\ :\ \vec{x}\in\mathbb{R}^{d-1},\tau\neq 0\Big{\}},$
$\mathrm{WF}\left(\big{(}x^{2}\big{)}_{-}^{-\frac{\nu}{2}}\right)=$$\Big{\{}\left((t,\vec{x}),(-\lambda t,\lambda\vec{x})\right)\ :\ t^{2}-\vec{x}^{2}=0,(t,\vec{x})\neq 0,\lambda\neq 0\Big{\}}$
$\cup\Big{\{}((0,0),(\tau,\vec{k}))\ :\ \tau^{2}-\vec{k}^{2}=0,(\tau,\vec{k})\neq 0\Big{\}},$
where $(\tau,\vec{k})$ denotes the dual variable to $(t,\vec{x})$. The fiberwise sum of wavefront sets $\mathrm{WF}\left(\theta(t)\right)+\mathrm{WF}\left(\left(x^{2}\right)_{-}^{-\frac{\nu}{2}}\right)$ does not contain an element of the form $((t,\vec{x}),(0,0))$. Therefore, by Hörmander’s criterion [20, p. 267], the product of these two distributions is well defined.
5.2 Green functions of the Klein Gordon equation
Consider the Klein-Gordon equation
$(-E-\Box)f(x)=g(x),$ (5.17)
where $E$ is a parameter, usually real. We will consider 3 cases:
massive case: $-E=m^{2},$ (5.18)
massless case: $-E=0,$ (5.19)
tachyonic case: $-E=-m^{2}.$ (5.20)
The massive and massless cases are quite similar and they often appear in physical applications. They are often discussed in detail in the literature. The tachyonic case is more exotic and less known, but also interesting.
The Klein-Gordon equation possesses several useful Green functions, that is, distributions satisfying
$(-E-\Box)G^{\bullet}(x)=\delta(x).$ (5.21)
[…]
$G_{m}^{\mathrm{F}/\overline{\mathrm{F}}}(x)=\int\frac{{\rm e}^{{\rm i}px}}{(-E+p^{2}\mp{\rm i}0)}\frac{{\rm d}p}{(2\pi)^{d}}.$ (5.22)
As follows from a general theory of hyperbolic equations, the Klein-Gordon equation (5.17) possesses also another important pair of Green functions: the retarded (or forward) Green function $G^{\vee}$ and the advanced (backward) Green function $G^{\wedge}$. They are uniquely defined by the conditions
${\rm supp}G^{\vee/\wedge}\subset J^{\pm}.$ (5.23)
Note that the above definition provides $G^{\vee/\wedge}$ for all $E\in{\mathbb{R}}$. In the case $-E\geq 0$, with $-E=m^{2}$ they coincide with $G_{m}^{\vee/\wedge}$ defined already with the help of Fourier transformation. In the tachyonic case they will be denoted $G_{{\rm i}m}^{\vee/\wedge}=G_{-{\rm i}m}^{\vee/\wedge}$ and they need a separate discussion, see Subsection 5.5
We will also consider certain distinguished solutions of the (homogeneous) Klein-Gordon equation, that is functions $G^{\circ}$ satisfying
$(-E-\Box)G^{\circ}(x)=0.$ (5.24)
One can look for them with the ansatz
$G^{\circ}(x)=\int{\rm e}^{{\rm i}px}g^{\circ}(p)\delta(-E+p^{2})\frac{{\rm d}p}{(2\pi)^{d-1}},$ (5.25)
where $g^{\circ}$ is a distribution on $p^{2}-E=0$. Above, for $E\in{\mathbb{R}}$, we use the notation
$\delta(p^{2}-E){\rm d}p=\frac{\delta\Big{(}p^{0}-\sqrt{\vec{p}^{2}-E}\Big{)}}{2\sqrt{\vec{p}^{2}-E}}{\rm d}\vec{p}+\frac{\delta\Big{(}p^{0}+\sqrt{\vec{p}^{2}-E}\Big{)}}{2\sqrt{\vec{p}^{2}-E}}{\rm d}\vec{p},$ (5.26)
where for $\vec{p}^{2}<E$ (5.26)=0.
Below we consider separately the massive, massless and tachyonic cases of the Klein-Gordon equation. In all three cases, we will be able to define $G^{\overline{\mathrm{F}}/\mathrm{F}}$ and $G^{\vee/\wedge}.$
5.3 Massive Klein-Gordon equation
Let us consider $-E=m^{2}$, that is the massive Klein-Gordon equation. The corresponding Green functions satisfy
$(m^{2}-\Box)G_{m}^{\bullet}(x)=\delta(x).$ (5.27)
Specifying Theorem 10 to $\mu=2$, we obtain the following expressions for the Feynman and anti-Feynman Green functions:
Theorem 12
$G_{m}^{\mathrm{F}/\overline{\mathrm{F}}}(x)=$ $\int\frac{{\rm e}^{{\rm i}px}}{(m^{2}+p^{2}\mp{\rm i}0)}\frac{{\rm d}p}{(2\pi)^{d}}$ (5.28)
$=$ ±$\frac{\pm{\rm i}\sqrt{\pi}m^{d-2}}{(4\pi)^{\frac{d}{2}}}U_{\frac{d}{2}-1}\Big{(}\frac{m^{2}(x^{2}\pm{\rm i}0)}{4}\Big{)}.$ (5.29)
The retarded and advanced Green functions of the Klein-Gordon equation are obtained by specifying Theorem 10 to $\mu=2$. In the following theorem, we also identify their regular and singular part.
Theorem 13
$G_{m}^{\vee/\wedge}(x)=$ $\int\frac{{\rm e}^{{\rm i}px}}{m^{2}+p^{2}\mp{\rm i}0{\rm sgn}p^{0}}\frac{{\rm d}p}{(2\pi)^{d}}$ (5.30)
$=$ $\theta(\pm x^{0})\frac{-{\rm i}\sqrt{\pi}m^{d-2}}{(4\pi)^{\frac{d}{2}}}\Bigg{(}U_{\frac{d}{2}-1}\Big{(}\frac{m^{2}x^{2}-{\rm i}0}{4}\Big{)}-U_{\frac{d}{2}-1}\Big{(}\frac{m^{2}x^{2}+{\rm i}0}{4}\Big{)}\Bigg{)}$ (5.31)
$=\theta(\pm x^{0})\frac{2\pi m^{d-2}}{(4\pi)^{\frac{d}{2}}}\big{(}\tfrac{x^{2}}{4}\big{)}_{-}^{1-\frac{d}{2}}{\bf F}_{1-\frac{d}{2}}\Big{(}\frac{m^{2}x^{2}}{4}\Big{)}.$ (5.32)
We can decompose $G^{\vee/\wedge}$ into a singular and regular part:
$G_{m}^{\vee/\wedge}(x)$ $=G_{m,{\rm sing}}^{\vee/\wedge}(x)+G^{\vee/\wedge}_{m,\mathrm{reg}}(x).$ (5.33)
For $d$ odd this decomposition can be chosen as
$G_{m,{\rm sing}}^{\vee/\wedge}(x)$ $=\frac{\theta(\pm x^{0})}{2\pi^{\frac{d}{2}-1}}\sum_{j=0}^{\frac{d-5}{2}}\frac{(-1)^{j}}{j!\Gamma(2-\frac{d}{2}+j)}\Big{(}\frac{m^{2}}{4}\Big{)}^{j}(x^{2})_{-}^{1-\frac{d}{2}+j},$ (5.34)
$G_{m,{\rm reg}}^{\vee/\wedge}(x)$ $=\frac{\theta(\pm x^{0})}{2\pi^{\frac{d}{2}-1}}\sum_{j=\frac{d-3}{2}}^{\infty}\frac{(-1)^{j}}{j!\Gamma(2-\frac{d}{2}+j)}\Big{(}\frac{m^{2}}{4}\Big{)}^{j}(-x^{2})^{1-\frac{d}{2}+j}\theta(-x^{2}).$ (5.35)
For $d$ even:
$G_{m,{\rm sing}}^{\vee/\wedge}(x)$ $=\theta(\pm x^{0})\frac{1}{2\pi^{\frac{d}{2}-1}}\sum_{j=0}^{\frac{d}{2}-2}\frac{(-1)^{j+1}}{(\frac{d}{2}-2-j)!}\biggl{(}\frac{m^{2}}{4}\biggr{)}^{\frac{d}{2}-2-j}\delta^{(j)}(x^{2}),$ (5.36)
$G_{m,{\rm reg}}^{\vee/\wedge}(x)$ $=\theta(\pm x^{0})\frac{2\pi m^{d-2}}{(4\pi)^{\frac{d}{2}}}{\bf F}_{\frac{d}{2}-1}\Big{(}\frac{m^{2}x^{2}}{4}\Big{)}\theta(-x^{2}).$ (5.37)
Proof.  The formula for the Green function of the Klein-Gordon equationis given by equation (5.32), which was computed earlier for a general $\mu$ (5.7). The decomposition into (5.34) and (5.35) is due to (2.48). For even $d$, the decomposition can be rewritten using (2.49). $\Box$
Introduce the following distinguished solutions of the Klein–Gordon equation $-\Box+m^{2}$:
$G_{m}^{\mathrm{PJ}}(x)$ $:=\frac{{\rm i}}{(2\pi)^{d}}\int{\rm e}^{{\rm i}x\cdot p}{\rm sgn}(p^{0})\delta(p^{2}+m^{2}){\rm d}p$ (5.38)
$=\frac{1}{(2\pi)^{d-1}}\int\frac{{\rm d}\vec{p}}{\sqrt{\vec{p}^{2}+m^{2}}}{\rm e}^{{\rm i}\vec{x}\vec{p}}\sin\left(x^{0}\sqrt{\vec{p}^{2}+m^{2}}\right)$ (5.39)
$G_{m}^{(\pm)}(x)$ $:=\frac{1}{(2\pi)^{d}}\int{\rm e}^{{\rm i}x\cdot p}\theta(\pm p^{0})\delta(p^{2}+m^{2}){\rm d}p$ (5.40)
$=\frac{1}{(2\pi)^{d-1}}\int\frac{{\rm d}\vec{p}}{2\sqrt{\vec{p}^{2}+m^{2}}}{\rm e}^{\mp{\rm i}x^{0}\sqrt{\vec{p}^{2}+m^{2}}+{\rm i}\vec{x}\vec{p}}.$ (5.41)
Following [13], we will call distinguished Green functions and solutions jointly propagators. $G_{m}^{\mathrm{PJ}}$ is supported in $J^{\vee}\cup J^{\wedge}$. Here are the expressions for these solutions in terms of positions:
$G_{m}^{\mathrm{PJ}}(x)=$ ${\rm sgn}(x^{0})\frac{2\pi}{(4\pi)^{\frac{d}{2}}}\Big{(}\frac{x^{2}}{4}\Big{)}_{-}^{\frac{2-d}{2}}{\bf F}_{\frac{2-d}{2}}\Big{(}\frac{m^{2}x^{2}}{4}\Big{)},$ (5.42)
$G_{m}^{(\pm)}(x)=$ $\frac{\sqrt{\pi}m^{d-\mu}}{(4\pi)^{\frac{d}{2}}}U_{\frac{d-2}{2}}\Big{(}\frac{m^{2}x^{2}\pm{\rm i}{\rm sgn}x^{0}0}{4}\Big{)}.$ (5.43)
Note the identities satisfied by the propagators:
$G_{m}^{\vee}-G_{m}^{\wedge}$ $=G_{m}^{\mathrm{PJ}}$ (5.44a)
$={\rm i}G_{m}^{(+)}-{\rm i}G_{m}^{(-)},$ (5.44b)
$G_{m}^{\mathrm{F}}-G_{m}^{\overline{\mathrm{F}}}$ $={\rm i}G_{m}^{(+)}+{\rm i}G_{m}^{(-)},$ (5.44c)
$G_{m}^{\mathrm{F}}+G_{m}^{\overline{\mathrm{F}}}$ $=G_{m}^{\vee}+G_{m}^{\wedge},$ (5.44d)
$G_{m}^{\mathrm{F}}$ $={\rm i}G_{m}^{(+)}+G_{m}^{\wedge}={\rm i}G_{m}^{(-)}+G_{m}^{\vee},$ (5.44e)
$G_{m}^{\overline{\mathrm{F}}}$ $=-{\rm i}G_{m}^{(+)}+G_{m}^{\vee}=-{\rm i}G_{m}^{(-)}+G_{m}^{\wedge}.$ (5.44f)
To prove these identities we use repeatedly
$\theta(\pm p^{0})2\pi{\rm i}\delta(p^{2}+m^{2})=\theta(\pm p^{0})\Big{(}\frac{1}{p^{2}+m^{2}-{\rm i}0}-\frac{1}{p^{2}+m^{2}+{\rm i}0}\Big{)},$ (5.45)
5.4 Massless Klein-Gordon equation
The massless case is quite similar to the massive one: we need only to set $m=0$ in the previous subsection. In particular, all identities (5.44) are satisfied. There are a few simplifications. Only the most singular part of the massive propagator remains in the massless case. This is the special case of Riesz potentials, massless limit of Bessel potentials, that we studied in the section 3.2
Theorem 14
$G_{0}^{\mathrm{F}/\overline{\mathrm{F}}}(x)=$$\pm\frac{{\rm i}\Gamma(\frac{d}{2}-1)}{4\pi^{\frac{d}{2}}}\big{(}x^{2}\pm{\rm i}0\big{)}^{1-\frac{d}{2}},$ (5.46)
$G_{0}^{\vee/\wedge}(x)=$ $\theta(\pm x^{0})\frac{1}{2\pi^{\frac{d}{2}-1}\Gamma(2-\frac{d}{2})}(x^{2})_{-}^{1-\frac{d}{2}},$ (5.47)
$G_{0}^{\mathrm{PJ}}(x)=$ ${\rm sgn}(x^{0})\frac{1}{2\pi^{\frac{d}{2}-1}\Gamma(2-\frac{d}{2})}(x^{2})_{-}^{1-\frac{d}{2}},$ (5.48)
$G^{(\pm)}_{0}(x)=$ $\frac{\Gamma(\frac{d}{2}-1)}{4\pi^{\frac{d}{2}}}\big{(}x^{2}\pm{\rm i}0{\rm sgn}(x^{0})\big{)}^{1-\frac{d}{2}}.$ (5.49)
For $d$ odd (5.47) and (5.48) can be rewritten as
$G^{\vee/\wedge}_{0}(x)$ $=\theta(\pm x^{0})\frac{(-1)^{\frac{d}{2}-2}}{2\pi^{\frac{d}{2}-1}}\delta^{(\frac{d}{2}-2)}(x^{2}),$ (5.50)
$G^{\mathrm{PJ}}_{0}(x)$ $={\rm sgn}(x^{0})\frac{(-1)^{\frac{d}{2}-2}}{2\pi^{\frac{d}{2}-1}}\delta^{(\frac{d}{2}-2)}(x^{2})$ (5.51)
Note that using (A.12) we can write identity (5.44d) as
$G^{\mathrm{F}}_{0}(x)+G^{\overline{\mathrm{F}}}_{0}(x)=G_{0}^{\vee}+G_{0}^{\wedge}=\frac{1}{2\pi^{\frac{d}{2}-1}}\rho_{-}^{\frac{d}{2}-1}(x),$ (5.52)
which agrees with the fact that massless retarded/advanced Green functions, also known as Riesz distributions (see [16]), are expressed by homogeneous distributions supported on $J^{\vee/\wedge}.$
5.5 Tachyonic Klein-Gordon equation
Let us now consider the tachyonic Klein-Gordon equation, which means, with $E=m^{2}$. Its Green functions satisfy
$(-m^{2}-\Box)G^{\bullet}(x)=\delta(x).$ (5.53)
Usually, tachyonic quantum fields are considered to be unphysical [21]. Nevertheless, every now and then there are attempts to analyze them in the physics literature, see [22], and more recently [23].
We have a minor notational problem how to indicate that we replaced $m^{2}$ with $-m^{2}$. Naively, one would think it should be indicated by both $+{\rm i}m$ and $-{\rm i}m$ instead of $m$. However, this would suggest the analytic continuation ${\rm e}^{{\rm i}\phi},$$\pm\phi\in[0,\pi]$, which is not always appropriate. This problem appears in the case of the Feynman propagator: we will write $G_{{\rm i}m}^{\mathrm{F}}$, but not $G_{-{\rm i}m}^{\mathrm{F}}$. Similarly, for the anti-Feynman propagator we will write $G_{-{\rm i}m}^{\overline{\mathrm{F}}}$, but not $G_{{\rm i}m}^{\overline{\mathrm{F}}}$. In the case of retarded/advanced propagators, this problem will be absent, since the analytic continuation can be performed in $m^{2}$: thus $G_{{\rm i}m}^{\wedge/\vee}=G_{-{\rm i}m}^{\wedge/\vee}$.
[…]
Theorem 15
$G_{{\rm i}m}^{\mathrm{F}}(x)/G_{-{\rm i}m}^{\overline{\mathrm{F}}}(x)$ $=\int\frac{{\rm e}^{{\rm i}px}}{(-m^{2}+p^{2}\mp{\rm i}0)}\frac{{\rm d}p}{(2\pi)^{d}}$ (5.54)
$=$ $\frac{\sqrt{\pi}m^{d-2}(\mp{\rm i})^{d+1}}{(4\pi)^{\frac{d}{2}}}U_{\frac{d}{2}-1}\Big{(}\frac{m^{2}(-x^{2}\mp{\rm i}0)}{4}\Big{)}.$ (5.55)
In particular, for $x^{2}>0$ we have
$G_{{\rm i}m}^{\mathrm{F}}(x)/G_{-{\rm i}m}^{\overline{\mathrm{F}}}(x)=\frac{\pm{\rm i}\sqrt{\pi}m^{d-2}{\rm e}^{\mp{\rm i}\pi(\frac{d}{2}-1)}}{(4\pi)^{\frac{d}{2}}}U_{\frac{d}{2}-1}\Big{(}\frac{m^{2}(-x^{2}\mp{\rm i}0)}{4}\Big{)}$
and for $x^{2}<0$
$G_{{\rm i}m}^{\mathrm{F}}(x)/G_{-{\rm i}m}^{\overline{\mathrm{F}}}(x)=\frac{\pm{\rm i}\sqrt{\pi}m^{d-2}{\rm e}^{\mp{\rm i}\pi(\frac{d}{2}-1)}}{(4\pi)^{\frac{d}{2}}}U_{\frac{d}{2}-1}\Big{(}\frac{m^{2}|x^{2}|}{4}\Big{)}.$
[…]
$G^{\mathrm{F}}_{{\rm i}m}(x)/G^{\overline{\mathrm{F}}}_{-{\rm i}m}(x)=$$\frac{\pm{\rm i}\sqrt{\pi}m^{d-2}{\rm e}^{\pm{\rm i}\pi(\frac{d}{2}-1)}}{(4\pi)^{\frac{d}{2}}}U_{\frac{d}{2}-1}\Big{(}\frac{{\rm e}^{\pm{\rm i}\pi}m^{2}(x^{2}\pm{\rm i}0)}{4}\Big{)}$ (5.56)
$=$ ±$\frac{\pm{\rm i}\sqrt{\pi}m^{d-2}{\rm e}^{\pm{\rm i}\pi(\frac{d}{2}-1)}}{(4\pi)^{\frac{d}{2}}}U_{\frac{d}{2}-1}\Big{(}\frac{m^{2}(-x^{2}\mp{\rm i}0)}{4}\Big{)},$ (5.57)
which coincides with (5.55). $\Box$
Unfortunately, the tachyonic Feynman and antiFeynman propagator do not have the usual physical interpretation, as the vacuum expectation value of the time-ordered, resp. anti-time-ordered product of fields. In fact, for tachyons the vacuum is ill defined. Nevertheless, some authors, e.g. [22], try to use the above Feynman propagator to define interacting tachyonic quantum field theory.
Retarded and advanced tachyonic Green functions $G_{{\rm i}m}^{\vee/\wedge}$ are not tempered distributions on ${\mathbb{R}}^{1,d-1}$, and therefore they cannot be expressed in terms of the Fourier transformation in all variables, as in the massive and massless cases (5.30). However, they are well-defined, and in the following theorem we give three equivalent formulas for these propagators.
Theorem 16
The forward and backward propagators in the tachyonic case are given by
$G_{{\rm i}m}^{\vee/\wedge}(x)=G_{-{\rm i}m}^{\vee/\wedge}(x)$
$=$ $\theta(-x^{2})\theta(\pm x^{0})\frac{\sqrt{\pi}m^{d-2}{\rm i}^{d+1}}{(4\pi)^{\frac{d}{2}}}\Bigg{(}U_{\frac{d}{2}-1}\Big{(}\frac{m^{2}(-x^{2})}{4}\Big{)}-U_{\frac{d}{2}-1}\Big{(}\frac{{\rm e}^{{\rm i}2\pi}m^{2}(-x^{2})}{4}\Big{)}\Bigg{)}$ (5.58)
$=$ $\theta(-x^{2})\theta(\pm x^{0})\frac{\sqrt{\pi}m^{d-2}(-{\rm i})^{d+1}}{(4\pi)^{\frac{d}{2}}}\Bigg{(}U_{\frac{d}{2}-1}\Big{(}\frac{{\rm e}^{-{\rm i}2\pi}m^{2}(-x^{2})}{4}\Big{)}-U_{\frac{d}{2}-1}\Big{(}\frac{m^{2}(-x^{2})}{4}\Big{)}\Bigg{)}$ (5.59)
$=$ $\theta(\pm x^{0})\frac{2\pi}{(4\pi)^{\frac{d}{2}}}\Big{(}\frac{x^{2}}{4}\Big{)}_{-}^{1-\frac{d}{2}}{\bf F}_{1-\frac{d}{2}}\Big{(}\frac{m^{2}|x^{2}|}{4}\Big{)},$ (5.60)
They are supported in $J^{\vee}$, resp. $J^{\wedge}$. We can decompose $G_{{\rm i}m}^{\vee/\wedge}$ into a singular and regular part:
$G_{{\rm i}m}^{\vee/\wedge}(x)$ $=G_{{\rm i}m,{\rm sing}}^{\vee/\wedge}(x)+G^{\vee/\wedge}_{{\rm i}m,\text{reg}}(x).$ (5.61)
For $d$ odd this decomposition is almost the same as (5.34), (5.35) but without the factor $(-1)^{j}$:
$G_{{\rm i}m,{\rm sing}}^{\vee/\wedge}(x)$ $=\frac{\theta(\pm x^{0})}{2\pi^{\frac{d}{2}-1}}\sum_{j=0}^{\frac{d-5}{2}}\frac{1}{j!\Gamma(2-\frac{d}{2}+j)}\Big{(}\frac{m^{2}}{4}\Big{)}^{j}(x^{2})_{-}^{1-\frac{d}{2}+j},$ (5.62)
$G_{{\rm i}m,{\rm reg}}^{\vee/\wedge}(x)$ $=\frac{\theta(\pm x^{0})}{2\pi^{\frac{d}{2}-1}}\sum_{j=\frac{d-3}{2}}^{\infty}\frac{1}{j!\Gamma(2-\frac{d}{2}+j)}\Big{(}\frac{m^{2}}{4}\Big{)}^{j}(-x^{2})^{1-\frac{d}{2}+j}\theta(-x^{2}).$ (5.63)
For $d$ even the decomposition is similar as in (5.36) and (5.37):
$G_{{\rm i}m,{\rm sing}}^{\vee/\wedge}(x)$ $=\theta(\pm x^{0})\frac{1}{2\pi^{\frac{d}{2}-1}}\sum_{j=0}^{\frac{d}{2}-2}\frac{1}{(\frac{d}{2}-2-j)!}\biggl{(}\frac{m^{2}}{4}\biggr{)}^{\frac{d}{2}-2-j}\delta^{(j)}(x^{2}),$ (5.64)
$G_{{\rm i}m,{\rm reg}}^{\vee/\wedge}(x)$ $=\theta(\pm x^{0})\frac{2\pi m^{d-2}}{(4\pi)^{\frac{d}{2}}}{\bf F}_{\frac{d}{2}-1}\Big{(}\frac{m^{2}|x^{2}|}{4}\Big{)}\theta(-x^{2}).$ (5.65)
Proof.  Our starting point is the formula (5.31) for the forward and backward propagator $G_{m}^{\vee/\wedge}(x)$. They are analytic in $m$. Therefore, we can apply the analytic continuation $m\mapsto{\rm e}^{{\rm i}\frac{\pi}{2}}m$:
$G_{{\rm e}^{{\rm i}\frac{\pi}{2}}m}^{\vee/\wedge}(x)=$ $\theta(\pm x^{0})\frac{-{\rm i}\sqrt{\pi}m^{d-2}{\rm e}^{{\rm i}\pi(\frac{d}{2}-1)}}{(4\pi)^{\frac{d}{2}}}\Bigg{(}U_{\frac{d}{2}-1}\Big{(}\frac{{\rm e}^{{\rm i}\pi}(m^{2}x^{2}-{\rm i}0)}{4}\Big{)}$
$\qquad-U_{\frac{d}{2}-1}\Big{(}\frac{{\rm e}^{{\rm i}\pi}(m^{2}x^{2}+{\rm i}0)}{4}\Big{)}\Bigg{)}.$ (5.66)
This yields (5.58). Alternatively, we can apply the analytic continuation $m\mapsto{\rm e}^{-{\rm i}\frac{\pi}{2}}m$, which yields (5.59). $\Box$
Let us compute the sum of the tachyonic Feynman and antiFeynman propagator:
$G_{{\rm i}m}^{\mathrm{F}}(x)+G_{-{\rm i}m}^{\overline{\mathrm{F}}}(x)$ =$=\begin{cases}\frac{2\pi m^{d-2}}{(4\pi)^{\frac{d}{2}}}{\bf F}_{\frac{d}{2}-1}\Big{(}-\frac{m^{2}x^{2}}{4}\Big{)},&\quad x^{2}>0;\\ \frac{4\pi}{(4\pi)^{\frac{d}{2}}}\Big{(}\frac{-x^{2}}{4}\Big{)}^{1-\frac{d}{2}}{\bf F}_{1-\frac{d}{2}}\Big{(}-\frac{m^{2}x^{2}}{4}\Big{)},&\quad x^{2}<0.\end{cases}$ (5.67)
Thus $G_{{\rm i}m}^{\mathrm{F}}(x)+G_{-{\rm i}m}^{\overline{\mathrm{F}}}(x)$ does not have a causal support, and consequently,
$G_{{\rm i}m}^{\mathrm{F}}(x)+G_{-{\rm i}m}^{\overline{\mathrm{F}}}(x)\neq G_{{\rm i}m}^{\vee}(x)+G_{{\rm i}m}^{\wedge}(x).$ (5.68)
The equality in (5.68) holds only for $x^{2}<0$.
Note that because of (5.68) we could not deduce the formulas of the forward and backward propagators from the Feynman and anti-Feynman propagators, and we had to apply a separate argument based on analytic continuation.
In the tachyonic case, we do not have the solutions $G_{{\rm i}m}^{(\pm)}$. However, we can define the Pauli-Jordan propagator
$G_{{\rm i}m}^{\mathrm{PJ}}(x)=G_{-{\rm i}m}^{\mathrm{PJ}}$ $=\frac{1}{(2\pi)^{d-1}}\int{\rm d}\vec{p}\,{\rm e}^{{\rm i}\vec{x}\vec{p}}\frac{\sin\left(x^{0}\sqrt{\vec{p}^{2}-m^{2}}\right)}{\sqrt{\vec{p}^{2}-m^{2}}}$ (5.69)
$={\rm sgn}(x^{0})\frac{2\pi}{(4\pi)^{\frac{d}{2}}}\Big{(}\frac{x^{2}}{4}\Big{)}_{-}^{1-\frac{d}{2}}{\bf F}_{1-\frac{d}{2}}\Big{(}\frac{m^{2}|x^{2}|}{4}\Big{)}.$ (5.70)
Note that $G_{{\rm i}m}^{\mathrm{PJ}}$ cannot be written in the form (5.25).
Among the identities (5.44) only (5.44a) is still true.
5.6 Averages of plane waves on the hyperbolic plane
The Minkowski space possesses two kinds of hyperboloids. The two-sheeted hyperboloid consists of two connected components isomorphic to the hyperbolic space. In this subsection, we compute the Fourier transform of the natural measure on one of the sheets of the two-sheeted hyperboloids, similarly as in the Theorem 4.
Consider the future/past hyperboloid in the $d$-dimensional Minkowski space, denoted ±${\mathbb{H}}_{\pm,m}={\mathbb{H}}_{\pm,m}^{d-1}$, consisting of points $p$ such that $p^{2}+m^{2}=0$ and ±$\pm p^{0}>0$. Let ${\rm d}\Omega_{m}$ denote the standard measure on ±,msubscriptplus-or-${\mathbb{H}}_{\pm,m}$. We will see that up to a coefficient its Fourier transform is essentially the “positive frequency solution of the Klein-Gordon equation.”
Theorem 17
±$\int_{{\mathbb{H}}_{\pm,m}}{\rm e}^{{\rm i}px}{\rm d}\Omega_{m}(p)=m^{d-1}\pi^{\frac{d-1}{2}}U_{\frac{d}{2}-1}\Big{(}\frac{m^{2}(x^{2}\pm{\rm i}{\rm sgn}x^{0}0)}{4}\Big{)}.$ (5.71)
Proof.  This average, up to a coefficient, coincides with $G_{m}^{(\pm)}$ defined in (5.40), which we have already computed:
±$\int_{{\mathbb{H}}_{\pm,m}}{\rm e}^{{\rm i}px}{\rm d}\Omega_{m}(p)$ $=2m\int{\rm e}^{{\rm i}px}\theta(\pm p^{0})\delta(p^{2}+m^{2}){\rm d}p$ (5.72)
$=(2\pi)^{d}mG_{m}^{(\pm)}(x).$ (5.73)
Therefore, it is enough to use the formula (5.43). $\Box$
5.7 Averages of plane waves on the deSitter space
The one-sheeted hyperboloid in the physics literature is usually called the deSitter space. It will be denoted $\mathrm{dS}_{m}=\mathrm{dS}_{m}^{d-1}$. It consists of points $p$ such that $p^{2}=m^{2}$. Let ${\rm d}\Omega_{m}$ denote the standard measure on $\mathrm{dS}_{m}$. We will compute the Fourier transform of the measure on $\mathrm{dS}_{m}$.
Theorem 18
$\int_{\mathrm{dS}_{m}}{\rm e}^{{\rm i}px}{\rm d}\Omega_{m}(p)=$
$=$ $m^{d-1}\pi^{\frac{d-1}{2}}\Bigg{(}{\rm i}^{d}U_{\frac{d}{2}-1}\Big{(}\frac{m^{2}(-x^{2}+{\rm i}0)}{4}\Big{)}+(-{\rm i})^{d}U_{\frac{d}{2}-1}\Big{(}\frac{m^{2}(-x^{2}-{\rm i}0)}{4}\Big{)}\Bigg{)}$ (5.74)
$=$ $\begin{cases}(-1)^{\frac{d}{2}}2m^{d-1}\pi^{\frac{d-1}{2}}\Bigg{(}U_{\frac{d}{2}-1}\Big{(}\frac{-m^{2}x^{2}\pm{\rm i}0}{4}\Big{)}\pm\sqrt{\pi}{\rm i}\big{(}\tfrac{-x^{2}}{4}\big{)}_{-}^{1-\frac{d}{2}}{\bf F}_{1-\frac{d}{2}}\Big{(}-\frac{m^{2}x^{2}}{4}\Big{)}\Bigg{)},&\frac{d}{2}\in{\mathbb{N}},\\ 2{\rm i}^{d-1}m^{d-1}\pi^{\frac{d}{2}}\big{(}\tfrac{-x^{2}}{4}\big{)}_{-}^{1-\frac{d}{2}}{\bf F}_{1-\frac{d}{2}}\Big{(}-\frac{m^{2}x^{2}}{4}\Big{)},&\frac{d}{2}\notin{\mathbb{N}}.\end{cases}$ (5.75)
Proof.
$\int_{\mathrm{dS}_{m}}{\rm e}^{{\rm i}px}{\rm d}\Omega_{m}(p)$ $=2m\int{\rm e}^{{\rm i}px}\delta(p^{2}-m^{2}){\rm d}p$ (5.76)
$=\frac{m}{\pi{\rm i}}\int{\rm e}^{{\rm i}px}\Big{(}\frac{1}{p^{2}-m^{2}-{\rm i}0}-\frac{1}{p^{2}-m^{2}+{\rm i}0}\Big{)}{\rm d}p$ (5.77)
$=\frac{m(2\pi)^{d}}{\pi{\rm i}}\big{(}G_{{\rm i}m}^{\mathrm{F}}(x)-G_{-{\rm i}m}^{\overline{\mathrm{F}}}(x)\big{)}.$ (5.78)
Then we can use the result for the tachyonic Feynman and anti-Feynman propagator (5.55). $\Box$
One can see that the singular part is different in even- and odd-dimensional cases.
Appendix A Appendix
A.1 Some identities
The following identities for $A>0$ follow from the 2nd Euler integral:
$\frac{1}{A^{\frac{\mu}{2}}}$ $=\frac{1}{\Gamma(\frac{\mu}{2})}\int_{0}^{\infty}{\rm e}^{-sA}s^{\frac{\mu}{2}-1}{\rm d}s,$ (A.1)
$\frac{1}{(A\pm{\rm i}0)^{\frac{\mu}{2}}}$ $=\frac{{\rm e}^{\mp{\rm i}\frac{\pi\mu}{4}}}{\Gamma(\frac{\mu}{2})}\int_{0}^{\infty}{\rm e}^{\pm{\rm i}tA}t^{\frac{\mu}{2}-1}{\rm d}t.$ (A.2)
We will also need the Fourier transform of the Gaussian function on the Euclidean space ${\mathbb{R}}^{d}$, and of the Fresnel function on the pseudo-Euclidean space ${\mathbb{R}}^{q,d-q}$ (with $q$ minuses):
$\int{\rm d}p{\rm e}^{-sp^{2}}{\rm e}^{{\rm i}px}$ $=\Big{(}\frac{\pi}{s}\Big{)}^{\frac{d}{2}}{\rm e}^{-\frac{x^{2}}{4s}},$ (A.3)
$\int{\rm d}p{\rm e}^{\pm{\rm i}tp^{2}}{\rm e}^{{\rm i}px}$ $=(\mp{\rm i})^{q}\Big{(}\frac{\pi}{t}\Big{)}^{\frac{d}{2}}{\rm e}^{\pm{\rm i}\frac{\pi}{4}d}{\rm e}^{\mp{\rm i}\frac{x^{2}}{4t}}.$ (A.4)
A.2 Distributions
In this paper, we often use the language of distributions on ${\mathbb{R}}^{d}$. We say that a distribution $T$ is regular if there exists a locally integrable function $f$ such that for a test function $\Phi$
$T(\Phi)=\int f(x)\Phi(x){\rm d}x.$ (A.5)
We will use the integral notation also for irregular distributions, e.g.
$\int\delta^{(n)}(x)\Phi(x){\rm d}x=(-1)^{n}\Phi^{(n)}(0).$ (A.6)
Let us now consider some special distributions on ${\mathbb{R}}$. For any $\lambda\in{\mathbb{C}}$
(±$(\pm{\rm i}x+0)^{\lambda}={\rm e}^{\pm{\rm i}\lambda\frac{\pi}{2}}(x\mp{\rm i}0)^{\lambda}:=\lim_{\epsilon\searrow 0}(\pm{\rm i}x+\epsilon)^{\lambda}.$
is a tempered distribution. If ${\rm Re}\lambda>-1$, then it is regular and given by the locally integrable function
${\rm e}^{\pm{\rm i}{\rm sgn}(x)\frac{\pi}{2}\lambda}|x|^{\lambda}.$ (A.7)
The functions
$x_{\pm}^{\lambda}:=|x|^{\lambda}\theta(\pm x)$ (A.8)
define regular distributions only for ${\rm Re}\lambda>-1$. We can extend them to $\lambda\in{\mathbb{C}}$ except for $\lambda=-1,-2,\dots$ by putting
$x_{\pm}^{\lambda}$ $:=\frac{1}{2{\rm i}\sin\pi\lambda}\Big{(}-{\rm e}^{-{\rm i}\frac{\pi}{2}\lambda}(\mp{\rm i}x+0)^{\lambda}+{\rm e}^{{\rm i}\frac{\pi}{2}\lambda}(\pm{\rm i}x+0)^{\lambda}\Big{)}.$ (A.9)
For $\lambda>-1$ (A.9) are regular and coincide with $\theta(\pm x)|x|^{\lambda}$. We have
$x_{\pm}^{\lambda+1}=|x|\cdot x_{\pm}^{\lambda}.$ (A.10)
Instead of $x_{\pm}^{\lambda}$, it is often more convenient to consider
$\rho_{\pm}^{\lambda}(x)$ $:=\frac{x_{\pm}^{\lambda}}{\Gamma(\lambda+1)}$ (A.11)
$=\frac{\Gamma(-\lambda)}{2\pi{\rm i}}\Big{(}{\rm e}^{-{\rm i}\frac{\pi}{2}\lambda}(\mp{\rm i}x+0)^{\lambda}-{\rm e}^{{\rm i}\frac{\pi}{2}\lambda}(\pm{\rm i}x+0)^{\lambda}\Big{)}.$ (A.12)
Note that using (A.11) and (A.12) we have defined $\rho_{\pm}^{\lambda}$ for all $\lambda\in{\mathbb{C}}$. We have
${\rm\partial}_{x}\rho_{\pm}^{\lambda}(x)=\pm\rho_{\pm}^{\lambda-1}(x).$
At integers we have
$\rho_{\pm}^{n}(x)$ $=\frac{x_{\pm}^{n}}{n!},\ \ \ n=0,1,\dots;$ (A.13)
$\rho_{\pm}^{-n-1}(x)$ =(±$=(\pm 1)^{n}\delta^{n}(x),\ \ \ n=0,1,\dots.$ (A.14)
Clearly, for ${\rm Re}(\lambda)\leq-1$ the distributions $\rho_{\pm}^{\lambda}$ are irregular.
Acknowledgement.
The support of the National Science Center of Poland under the grant UMO-2019/35/B/ST1/01651 is acknowledged. We thank the referees for their remarks. In particular, we are grateful to one of them for drawing our attention to Hörmander’s criterion for multiplication of distribution; see its application at the end of Subsection 5.1